### v1

# Transportability of Prognostic Markers: Rethinking Common Practices through a Sufficient-Component-Cause Perspective

Abstract
Transportability, the ability to maintain performance across populations, is a desirable property of of markers of clinical outcomes. However, empirical findings indicate that markers often exhibit varying performances across populations. For prognostic markers whose results are used to quantify of the risk of an outcome, oftentimes a form of updating is required when the marker is transported to populations with different disease prevalences. Here, we revisit transportability of prognostic markers through the lens of the foundational framework of sufficient component causes (SCC). We argue that transporting a marker “as is” implicitly assumes predictive values are transportable, whereas conventional prevalence-adjustment shifts the locus of transportability to accuracy metrics (sensitivity and specificity). Using a minimalist SCC framework that decomposes risk prediction into its causal constituents, we show that both approaches rely on strong assumptions about the stability of cause distributions across populations. A SCC framework instead invites making transparent assumptions about how different causes vary across populations, leading to different transportation methods. For example, in the absence of any external information other than disease prevalence, a cause-neutral perspective can assume all causes are responsible for change in prevalence, leading to a new form of marker transportation. Numerical experiments demonstrate that different transportability assumptions lead to varying degrees of information loss, depending on how population differ from each other in the distribution of causes. A SCC perspective challenges common assumptions and practices for marker transportability, and proposes transportation algorithms that reflect our knowledge or assumptions about how causes vary across populations.
footnotetext: From Faculty of Pharmaceutical Sciences and Faculty of Medicine (MS), Vancouver, Canada; School of Population Health, WHO Collaborating Centre for Climate Change and Health Impact Assessment, and enAble Institute, Faculty of Health Sciences, Curtin University, Perth, Australia (GP); Faculty of Medicine, Universitas Negeri Malang, Indonesia (GP); Saw Swee Hock School of Public Health, National University of Singapore, Singapore (WC)footnotetext: * Correspondence to Mohsen Sadatsafavi, Room 4110, Faculty of Pharmaceutical Sciences, 2405 Wesbrook Mall, Vancouver, British Columbia, V6T1Z3, Canada; email: mohsen.sadatsafavi@ubc.ca
Background
The underlying premise of reporting on the performance of biomarkers, tests, and prediction models (which we generally refer to as “markers”) is that performance metrics are assumed to be transportable from one population to another. In this context, we define transportability as conditions under which certain indices of marker performance remain the same in the new population as in the source population (where it was developed). Examples of such indices include accuracy (sensitivity and specificity), predictive values, discrimination, and calibration. However, in practice, this premise more often than not fails to hold. Research has repeatedly shown the performance of markers can vary significantly across populations, especially when the prevalence of the outcome of interest differs1. For example, risk prediction models for cardiovascular disease have shown substantial degradation in performance when transported to a new population, leading to substantial risk of harm (defined as the net benefit of using the model being lower than not using it)2. The likelihood of harm was reduced when models were updated to account for the difference in prevalences between their source population and the target population. A similar phenomenon was observed in risk prediction models for acute exacerbations of obstructive lung disease3. In another example, the performance of machine learning models for detecting pneumonia on chest X-rays substantially declined in data from settings not used to train the model4.
Often times, when transporting a marker for an outcome to a new population, some information about the characteristics of that population - colloquially referred to as ‘case-mix’- is available. One of the most common information about the new population is the outcome prevalence. For example, cancer registries often provide a good estimate of cancer risk in a population. This poses the challenge as to how such information can be used to revise our assessment of marker performance. For risk predictions models that return a quantitative estimate of disease (or outcome) risk given patient characteristics, some form of model updating (or revision) is often needed to correct for the under-estimation or over-estimation of risks in the new population5. When the model is a logistic regression, the most basic form of such updating involves modifying the intercept of the model to account for difference in prevalence between the source and target populations1,6,7. The underlying justification for this approach is that differences in outcome prevalence are mostly due to factors that act independently of predictors used in the model6. Because in a logistic regression model, changing the intercept is equivalent to applying a correcting odds-ratio to predicted risks, this method can be generalized to correcting the outputs of any algorithm that predicts the risk of a binary outcome, including black-box (e.g., machine learning) models8.
The transportability of predictive information across populations is an active area of research in predictive analytics and machine learning. A recent scoping review categorized methods aimed at developing transportable marker, or making an existing marker transportable to a new setting, based on whether they require access to data from the target population, and whether they are purely data-driven or require contextual knowledge about associations9. One common underlying framework in knowledge-driven approaches is causal graphs, particularly the pioneering work on selection diagrams by Pearl et al.10. Generally, these methods aim to identify and remove predictors whose association with the outcome varies across populations. More recent developments include the creation of graph surgery estimators that break the link between selection into the population and causal associations between the outcome and predictors whose association with the outcome is not consistent11,12. This approach has recently been used to study how common metrics of model performance for prognostic and diagnostic markers change across populations with difference case-mix13.
Causal graphs are not the only model for causation. Another is the sufficient component causes (SCC) model14,15. SCC is a foundational model in that it establishes fully deterministic relationships between causes and effects (rather than a representation of statistical dependencies as in causal graphs). In this framework the variability in the outcome is due to the variability of, and interactions between, causes. This framework has recently been applied to explore the biologic plausibility of different link functions for modeling binary outcomes, resulting in proposals for more transportable measures of association such as the ‘switch relative risk’16 and more biology-aligned statistical models such as regression by composition17.
In this paper, we use a simple SCC model to study the most basic prediction setup: a binary factor that is used as a marker for the risk of an outcome. As a reference, we formulate this setup for prognostic markers, where adjustment for outcome prevalence seems to be a topical issue. We discuss how logical symmetries can be used to arrive at interpretations applicable to diagnostic markers. We create minimalist logic circuits to model a prognostic marker that provides partial, inconclusive information about an outcome of interest. This setup is the used to study marker transportability across populations, particularly as it relates to the variability in outcome prevalence. Our thesis is that by reducing the transportability problem to its basic constituents in this model, patterns will emerge that can provide insight into more complex scenarios.
A parsimonious SCC framework for prognostic markers
The SCC framework assumes the existence of sets of sufficient causes that bring about an event18. Within each set, the causes are non-redundant (all elements with the set are required for the event to happen), but sets can act independently of each other19.
Consider a binary prognostic marker, such as the presence or absence of BRCA1 mutation, and a binary outcome, such as the occurrence of breast cancer. While BRCA1 mutation is associated with increased cancer risk20, it is not a definitive marker: neither a negative BRCA1 mutation eliminates the risk of breast cancer, nor does its presence guarantee that cancer will occur. The fact that neither positive nor negative marker values are definitive indicates that there are at least two other mechanisms at play. On one hand, a BRCA1-positive individual must experience some other, key events that ultimately lead to cancer, explaining why not everyone with a BRCA1 mutation will develop cancer. On the other hand, cancer can also occur via pathways independent of the BRCA1 mutation, explaining why some breast cancer cases are BRCA1-negative.
We now formalize a minimal setup that explains such indeterminism of a marker value for an outcome. We are interested in the properties of a binary prognostic marker $T$ for a binary outcome $D$. For this marker to be informative but not definitive, at least two latent variables (or switches) must be present that can cause false-negative and false-positive responses. We can model these switches as follows:
  • The latent binary variable $U$ represents universally required causes - for example factors that cause pre-cancers (in-situ cancer) to progress to malignancy. The absence of $U$ is responsible for false-positive marker values.
  • The latent binary variable $V$ represents all alternative causes - for example pathways related to the effect of tobacco smoking, which increases the risk of breast cancer even among those without a BRCA1 mutation. The absence of $V$ is responsible for false-negative marker values.
We assign the values of 1 and 0, respectively, to the ‘on’ and ‘off’ status of each of these switches. For the brevity of notations, by a simple character, we refer to the ‘on’ value, and by its dot-accented to its negated value (e.g., $\dot{U}=\neg U$ indicating $U=0$). We use the shorthand notation $P(X)$ for $P(X=1)$. For example, $P(T\dot{V})=P(T\dot{V}=1)=P(T=1,V=0)$.
The above setup results in the following outcome-generating process for prognostic markers:
$D=(T\land U)\lor(V\land U)=(T\lor V)\land U.$
Before proceeding, we note that one can create an equally parsimonious SCC setup by swapping the logical AND and OR in the above equation, resulting in the setup $D=(T\land V)\lor U$. However, this is mathematically symmetrical to our reference setup. This becomes obvious by considering that the complementary marker (whose positive and negative results are swapped) is a marker for not experiencing the outcome under the reference setup: $\dot{D}=(\dot{T}\lor\dot{V})\land\dot{U}$. Because of this, any pattern we observe for PPV is also observed for NPV in this alternative formulation, albeit in the opposite direction, and vice versa. Similarly, any pattern we observe for SE in the reference setup is also observable, in the opposite direction, for SP in this alternative setup. Table 1 shows the minimum SCC setup, and the resulting 2-by-2 contingency tables of marker and outcome values ($P(T,D)$), from which performance metrics can be derived.
Table 1: Minimal configurations for a prognostic for the reference (left) and its symmetrical (right) setup, and the resulting contingency tables. Top: logic equation; Middle: SCC diagram; Bottom: The resulting contingency table
Reference setup Reciprocal setup
$D=(T\lor V)\land U$ $D=(T\land V)\lor U$
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Given that $T$, $U$, and $V$ are all binary, a population is made up of 8 subgroups. For example, $T\dot{U}V$ is the subgroup where $T=1$, $U=0$, and $V=1$. As $U$ and $V$ are latent variables, the observed properties of the marker in a population is manifested in terms of $P(T,D)$, i.e., the two-by-two (contingency) table of marker by outcome status probabilities in that population.We express a contingency table by the sequence TP,FP,FN,TN{TP,FP,FN,TN}, where $TP=P(TD),FP=P(T\dot{D}),FN=P(\dot{T}D),TN=P(\dot{T}\dot{D})$.
If all we know about populations is their contingency tables, the full joint distribution $P(T,U,V)$ is unidentifiable, given its specification requires seven degrees of freedom but the contingency table only provides three. However, any joint distribution can be replaced by independent Bernoulli distributions for $P(T)$, $P(U)$, and $P(V)$ that would result in the same contingency table. Given specifying three Bernoulli distributions requires three degrees of freedom, this setup is identifiable and has a unique (1:1) mapping to a given contingency table. Because our focus is on metrics that are determined by the contingency table, without loss of generality the following we assume independence i.e., $P(T,U,V)=P(T)P(U)P(V)$.
Our interest is in the four fundamental performance characteristics of binary markers: sensitivity (SE=P(T|D)=TP/(TP+FN){SE}=P(T|D)=TP/(TP+FN)), specificity (${SP}=P(\dot{T}|\dot{D})=TN/(FP+TN)$), positive predictive value (PPV=P(D|T)=TP/(TP+FP){PPV}=P(D|T)=TP/(TP+FP)), and negative predictive value (${NPV}=P(\dot{T}|\dot{D})=TN/(FN+TN)$), as a function of prevalence ($P(D)=TP+FN$). By convention by SE and SP we refer to accuracy metrics, and by PPV and NPV as predictive values.
As a numerical example, consider a population where $P(T)=$ 0.25, $P(U)=$ 0.75, and $P(V)=$ 0.50. The 8 subgroups are scattered into the four cells of the contingency table (see Table 1). The contingency table emerging from this setup is {0.188,0.062,0.281,0.469}.This results in a prevalence ($P(D)$) of 0.469, SE of 0.400, SP of 0.882, PPV of 0.750, and NPV of 0.625. We will use this simple numerical example as a case study on various ways a marker can be transported.
Two common methods of transportability
In this section we review the two conventional approaches for transporting a marker to a new population with potentially different outcome prevalence: transportation via accuracy (SE & SP), and transportation via predictive values (PPV & NPV).
In transportation by accuracy, we assume the SE and SP are equal between the source and target populations. The contingency table in the target population can be recovered by knowing outcome prevalence in the target population, denoted by $\pi$. For example, the proportion of true positives consists of those with the disease who also rest positive: $TP=\pi\times SE$, and so on. From this contingency table, various performance metrics can be derived. Details of this approach are provided in Appendix 1 (section A). This approach is the modus operandi for binary diagnostic markers, where the Bayes’ rule is used to combine pre-test probability with marker accuracy estimates to derive the post-test probability of the outcome21,22. Without any information that would distinguish the individual under evaluation, the pre-test probability is taken to be disease prevalence in the target population, which is equal to this mode of transportation.
Transportation by predictive values refers to the where we assume it is the PPV and NPV from the source population that are transportable. Under this assumption, differences in outcome prevalence is due to the change in the prevalence of marker-positive individuals across populations. Again, if the target prevalence is known, PPV and NPV can be used to recover the contingency table in the target population. Details of this approach are provided in Appendix 1 (section B). This approach is more common for prognostic markers, a typical example being risk scoring tools that directly return an estimate of $P(D|T=x)$ for someone with marker value $x$. For a binary marker, this risk equation can be written as P(D|T=x)=(1NPV)+[PPV(1NPV)]xP(D|T=x)=(1-{NPV})+[{PPV}-(1-{NPV})]x. This equation indicates the implied assumption that the predictive values of the marker are transportable. In our numerical example, advertising prognostic information as a risk equation would result in $P(D|T)=0.375+0.375T$.
Returning to our numerical example, consider a population with 25% higher prevalence of breast cancer ($P(D)=0.586$). Transportation by accuracy results in a contingency table of {0.234,0.049,0.352,0.365}, corresponding to a PPV of 0.828 and NPV of 0.510. On the other hand, transportation by predictive values gives rise to a contingency table of {0.422,0.141,0.164,0.273}, corresponding to a SE of 0.720 and SP of 0.660.
Prevalance adjustment
A common way that prognostic markers are transported across populations is in the format of a risk equation for $P(D|x)$, either explicitly, as is the case in regression-based prediction models, or implicitly, as in black-box (e.g., machine learning) models. Consider applying the Bayes’ rule to the outcome risk conditional on marker value:
$P(D|T=x)=\frac{P(D)P(T=x|D)}{P(x)}=\frac{P(D)P(T=x|D)}{P(D)P(x|D)+P(\dot{D})P(x|\dot{D})}.$
A familiar modeling framework for binary outcomes is the logistic regression. There, the logit function ($\mbox{logit}(x)=\log(\frac{x}{1-x})$) is used as link function connecting marker value to outcome probability. Applying this function to both sides of the risk equation for marker values, we have
$\mbox{logit}(P(D|T=x))=\mbox{logit}(P(D))+\log\left(\frac{P(T=x|D=1)}{P(T=x|D=0)}\right).$
The last term on the right-hand side is the likelihood ratio (LR) of the marker between the diseased and non-diseased groups. The first term on the right-hand side is the logit of prevalence, which is not a function of marker value. These derivations indicate that the practice of prevalence adjustment by the odds-ratio of prevalence between the source and target populations is equivalent to the application of Bayes’ theorem (where outcome prevalence is taken as the prior probability). This approach changes the locus of transportability from predictive values to the LR: for prevalence adjustment to make the marker transportable, the LR must remain constant across all levels of the marker.
For binary markers, the LR is defined at two values. For $T=0$ it is (1SE)/SP(1-SE)/SP (aka negative LR), and for $T=1$ it is SE/(1SP)SE/(1-SP) (aka positive LR). For the marker to be transportable, both these LRs need to remain constant across populations, which will be the case if and only if both SE and SP remain constant. This in turn implies that prevalence adjustment for binary markers is equal to considering SE and SP to be transportable.
Under what conidtions are performance metrics transportable?
The above reasoning shows why adjustment for prevalence is generally expected to improve the transportability of a marker, as the conventional wisdom is that SE and SP, being defined within the diseased and non-diseased groups, are less dependent on prevalence23. While this might be intuitive as a general observation, under the SCC framework, none of these metrics are truly intrinsic. Rather, they emerge as properties of the distribution of causes, and their transportability depends on how this distribution varies across populations. One can indeed create population-generating mechanisms where a given subset of these metrics remains constant while others vary. However, the SCC framework helps us examine the plausibility of such mechanisms.
Table 2 shows conditions under which each of the four metrics remains constant (and therefore transportable) for a prognostic marker.
Table 2: Conditions for Transportability of Metrics of Marker Performance
Metric Transportability condition
SE{SE} ${P(\dot{T}UV)}/{P(TU)}=c$
SP{SP} ${P(T\dot{U})}/({P(\dot{T})-P(\dot{T}UV)})=c$
PPV{PPV} $P(U\mid T)=c$
NPV{NPV} $P(UV\mid\dot{T})=c$
SE: sensitivity; SP: specificity; PPV: positive predictive value; NPV: negative predictive value
  • As an example of derivations, consider PPV. Its transportability means $P(D|T)=c$ (a constant). But $P(D|T)=(P(TU\dot{V})+P(TUV))/P(T)=P(TU)/P(T)=P(U|T)$.
For the PPV to be transportable, the prevalence of universal causes among marker-positive individuals should remain constant. For NPV, transportability requires that among test-negative individuals, the proportion in whom both the universal and alternative causes are present should remain constant. These conditions are both satisfied if the prevalences of $U$ and $V$ are stable across populations. In this scenario, the entirety of variation in outcome prevalence is attributable to variation in $T$. In our breast cancer example, this would mean populations vary in breast cancer prevalence only because they differ in the BRCA1 mutation rate. For many markers, including BRCA1, this assumption is unrealistic, as many factors contribute to the variability in breast cancer risk24. In comparison, conditions for the transportation by accuracy are more complicated. It does not seem possible to specify a simple model on how $P(T,U,V)$ should vary such that both SE and SP remain stable. Such transportability method for a prognostic marker, as is implied in prevalence adjustment, amounts to a very specific set of assumptions on how causes vary across populations.
How do performance metrics vary by prevalence under different population-generating mechanisms?
In the SCC framework, the degree of transportability of a marker by a given performance metric depends on how the causes vary across populations. Despite the simple setup, complex patterns arise due to the non-linearities in the interplay among causes. In this section we visualize how the relationship between performance metrics and outcome prevalence changes under various population-generating mechanisms. Taking the population of our numerical example as the baseline, we modeled the following scenarios: when populations vary only in one of causes; when they vary in two of the three causes; and when they vary in all three causes. When multiple causes vary, we assume they change by the same degree on the odds-ratio scale.
Results are provided in Figure 1. As explained above, under $T$-only variation, PPV and NPV remain transportable. This again highlights that transporting a marker by its predictive values (as in risk equations) means attributing change in prevalence entirely due to $T$, a very strong assumption in most real scenarios. Importantly, in all other conditions, PPV and NPV varied by prevalence. On the other hand, in none of the modeled scenarios did SE and SP remain unchanged. In fact, SE and SP could vary in either direction as a function of prevalence, reflecting the more complex conditions required for their transportability.
Figure 1: Relationship between prevalence and marker performance metrics under various population-generating mechanisms for a prognostic marker
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
SE: sensitivity; SP: specificity; PPV: positive predictove value; NPV: negative predictive value
  • The base setup is $P(T)=0.25$, $P(U)=0.75$, $P(V)=0.5$. Changes across populations are modeled by applying varying odds-ratios to the components that vary (labeled on top of each panel).
Other forms of marker transportation under the SCC framework
The SSC framework enables us to express our beliefs on how causal pathways vary across populations, resulting in algorithms that are transparent and explicit in their underlying assumptions. Consider investigating a rare ancestral mutation as a prognostic marker for cancer that is known to have minimal variations across populations. In this case, it is reasonable to assume that $P(T)$ remains essentially constant across populations. When transporting this marker, our focus would therefore be on scenarios where differences in $P(U,V)$ derive variations in cancer prevalence. The resulting transportation method will be different from transportation by predictive values or accuracy. Indeed, both these transportation methods implicitly require that an increase in outcome prevalence be accompanied by change in marker positivity.
But, what if the only information we have is differences in outcome prevalence across populations? We believe it is still worth visiting the plausibility of the assumptions under various transportability methods even under such a general case. For transportability by predictive values, one can question as to why only one cause ($T$, the one we happen to be measuring) should be responsible for change in prevalence. Transportability by SE and SP is also questionable due to unclear assumptions it places on the distribution of causes. For this general case, it might be more reasonable to take an explicit yet neutral stance about the variations in causes that bring about variations in prevalence, for example that all causes move by the same extent to cause change in prevalence - this approach makes no exception for thae causes for the mere fact that we are measuring it. As an implementation of this approach, one can solve for a common odds-ratio that, when applied to $P(T)$, $P(U)$, and $P(V)$ in the source population, results in a new population that matches the outcome prevalence in the target population. This “proportional odds” assumption is a new, distinct transportability approach that would generate different updated values of marker performance metrics compared with both conventional methods. Details of operationalizing this approach are provided in Appendix 1 (section C).
In our numerical example, transporting the marker to a new population with 25% higher prevalence would require applying an odds-ratio of 1.434 to the three causes. This odd-ratio can be derived via standard univariate root finding (given that there is a monotonical relationship between this odds-ratio and prevalence). It can also be expressed as the real root of a cubic equation (see Appendix 1 - section C). The contingency table under this method of transportability will be {0.262,0.061,0.323,0.353}. This in turn results in SE of 0.448 and SP of 0.853. The PPV and NPV are, respectively, 0.811 and 0.522. Formulating marker performance as a risk equation would thus result in $P(D|T)=0.189+0.333T$). Note that both the intercept and slope (coefficient for $T$) of the risk equation are updated, which is different from the risk equation under the assumption of transportability of predictive values (where both remain unchanged), or accuracy (where only the intercept is updated).
Information loss under different transportability assumptions
Ultimately, any assumption about how causes vary across populations will be a simplified version of a complex reality. The discrepancy between marker performance under a given transportability assumption and the true marker performance results in loss of prognostic information. We conducted brief simulation studies to explore such information loss under simple population-generating scenarios. The discrepancy between an assumed versus the true marker performance can be measured in different ways, each focusing on different consequences of such discrepancy (for example, through measures of discrimination, calibration, or prediction error). A more foundational approach is to measure information loss according to information theory, which in this context can be measured via the Kullback-Leibler divergence ($D_{KL}$)25. This information-theoretic measure quantifies the discrepancy between a true distribution and a candidate distribution in terms of ‘excess surprise’, which can be interpreted as the additional number of bits required to encode information from the true distribution using the candidate distribution, rather than using the true distribution itself26. In our case, these distributions are the true contingency table ($P(T,D)$) in the target population versus the one implied by a given transportability algorithm.
We modeled population-generating scenarios similar to those presented in Figure 1. Because in realistic settings all three causes are likely to vary, we explored this setup with more depth. We modeled scenarios where changes in the three causes had different strengths of correlation with each other (ranging from changes being uncorrelated to fully correlated). We also modeled a ‘maximum entropy’ scenario where the prevalence of the eight subgroups follows Dirichlet(1,1,1,1,1,1,1,1); this distribution has maximum information-theoretic randomness (entropy)27. For each scenario, we simulated random pairs of source and target populations according to a given population-generating mechanism. For each pair, we applied the following three transportation strategies: by accuracy (SE and SP), by predictive values (PPV and NPV), and via the proportional-odds assumption explained previously. Divergence metrics were averaged across 10,000 simulations for each scenario.
Results are provided in Figure 2. This confirms that if the outcome prevalence variation is entirely due to $T$, PPV and NPV are fully transportable ($D_{KL}=0$). However, under other scenarios, transporting by predictive values resulted in substantial information loss. Transporting by accuracy reduced information loss in some scenarios, albeit in none was the loss zero, indicating that none of the modeled scenarios were compatible with stable SE and SP. Transporting by accuracy performed worse than transporting by predictive values in some scenarios, indicating that prevalence adjustment is not guaranteed to improve the transportability of a prognostic marker. The proportional-odds assumption was also reduced information loss in some scenarios, but increased it in others. For the scenarios where the three causes changed together, transportation by accuracy generally reduced information loss compared with transportation by predictive values (unless the correlation among causes were weak). In all those situations, transportation via he proportional-odds assumption performed better than the conventional transportability methods. As expected, in the scenarios where changes in causes were positively correlated.
Figure 2: Information loss associated with different methods for transporting a marker under various population-generating scenarios
Refer to caption Refer to caption
  • Kullback-Leibler divergence $(D_{KL})=\sum_{i=1}^{4}P(a_{i})\log_{2}(P(b_{i})/P(a_{i}))$ with $a_{1},a_{2},a_{3},a_{4}$ being the four cells of the contingency table from the correct $P(T,D)$, and $b_{i}$s the corresponding ones from $P(T,D)$ constructed given the transportability method.
  • The base population is as follows: $P(T)=0.25$, $P(U)=0.75$, and $P(V)=0.5$. When a cause varies, it is given a logit-normal distribution with mean being equal to the base population’s point estimate, and standard deviaioon of 0.1.
  • No coorelation, weak, moderate, strong, and full correlation correspond to correlation coefficients of, respectively, 0, 0.25, 0.50, 0.75, amd 1 for logit-transformed probabilities.
Discussion
We constructed a simple causal model under the SCC framework to study the transportability of prognostic markers that are partially informative about an outcome of interest but also generate false-positive and false-negative results. Several observations from this theoretical exploration deserve being highlighted. First, the common practice of transporting the marker by predictive values is questionable, as it relies on the strict assumption that the prevalences of causes other than those represented by the marker remain stable across populations. Further, we showed that the conventional prevalence adjustment method is the exact implementation of Bayes’ rule, thus changing the locus of transportability from predictive values to accuracy metrics (sensitivity and specificity). Nevertheless, no easily explainable or biologically plausible mechanism is likely to generate fully transportable accuracy metrics. In our explored scenarios regarding the distribution of causes across populations, transporting a prognostic marker by accuracy was not universally better than transporting by predictive values. However, in scenarios where causes were positively correlated, which is likely to be common, transporting by accuracy did reduce information loss compared with transportability by predictive values.
These results are based on a selected set of population-generating mechanisms and a single set of core parameters for the SCC model ($P(T)=0.25$, $P(U)=0.75$, $P(V)=0.5$). Still, they should be sufficient to question some common assumptions and practices. These include the practice of advertising prognostic risk questions without emphasizing their dependence on prevalence, or considering sensitivity and specificity as intrinsic properties of markers. While these insights are from studying binary markers, the core findings can be extended to continuous markers and multi-variable risk equations. Transforming such markers ‘as is’ is equal to assuming that stratum-specific predictive values ($P(D|T=x)$) remain constant. This is equal to attributing variations in prevalence entirely to variations in $P(T)$ (aka case-mix). Consider a multi-variable risk score such as the QRISK for cardiovascular diseases28. Even though this model includes up to 22 predictors, is it plausible that other causal pathways to cardiovascular diseases that might vary across populations, including yet-to-be-discovered genetic risk factors, environmental exposures, subtle lifestyle choices, access to and quality of preventive care, and so on, are the same across populations? This assumption is needed for claiming that QRISK predictions, developed using primary care UK data, are transportable to other settings. On the other hand, conventional prevalence adjustment would indicate that stratum-specific likelihood ratio, i.e., $P(x|D)/P(x|\dot{D})$ would be transportable. This assumption will hold if the distribution of predicted risks among those who will experience the outcome and those who will not remains is same across diverse populations. This condition places very specific constraints on how the causes of diseases should vary across populations.
We focused on prognostic markers given the recent debates on transportabiltiy of prognostic information and the merit of prevalence-adjustment to improve transportability. The application of SCC framework for diagnostic markers is also important and deserves its own airing. For such markers, the path of causality if from the disease to the test. As such, an equivalent setup to our reference setup for a diagnostic marker would be $T=(D\lor V)\land U$. The conventional method of transporting diagnostic markers is by accuracy (e.g., the Bayes’ rule, which combines accuracy metrics from the source population with disease prevalence from the new population). From this setup, it is immediately obvious that transportation by accuracy requires that universal and alternative causes $U$ and $V$ remain stable across populations. This assumption is similar to the assumption of transportability by predictive values for prognostic markers. This strong requirement can explain why these metrics tend to vary across populations29–31. Further details of how the SCC framework can inform transportability of diagnostic markers needs to be pursued in future work.
How can these findings inform practice? If reliable, context-specific information is available on the distribution of specific predictors and their relationships with the outcome and with each other, methods based on causal graphs can be used to build transportable models or to design tailored transportation strategies11. In contrast, when only general information, such as outcome prevalence or test positivity rates are available, the SCC framework can be used to formulate tailored transportability algorithms that utilize such information. We question the contemporary practice of advertising risk prediction models for outcome risk as transportable. We believe risk equations should explicitly mention the outcome risk in their source population, and prevalence adjustment should be advocated as a necessary step before using the equation in a population with different outcome prevalence. Yet, while conventional prevalence adjustment may serve as a reasonable default, it cannot be assumed to consistently improve performance. Instead of fixed transportability rules applied to common performance metrics, the SCC framework offers a foundation for updating marker performance in ways that are more grounded in mechanistic interactions among potential causes. This results in new transportability algorithms based on assumptions on how underlying causes vary, rather than keeping performance metrics, which are emergent properties of such causes, constant. An example is the proportional-odds prevalence adjustment method discussed above. It is motivated by the notion that oftentimes differences in outcome prevalence are multi-factorial. An agnostic stance that that treats all causes as equally responsible is more defendable than blaming only one (as in transportation by predictive values) or assuming they move in unxplainable ways just to keep sensitivity and specificity constant (as in transportability by accuracy). Of course, variations in this method are also possible. Imagine, for example, that in addition to knowing target prevalence, we have knowledge of marker positivity rate ($P(T)$) in the new population (e.g., from an independent cross-sectional study). One can then directly adjust $P(T)$ to match this value, and then apply a common odds-ratio to other causes to match target prevalence.
We conclude this paper by providing a few areas for further inquiry. Whether SCC-informed transportation algorithms such as the proportional-odds method actually improve transportability can be tested in empirical studies. Our arguments were based on precise knowledge of outcome prevalence in the target population. To what extent using a noisy estimate of prevalence will help or harm transportability needs to be investigated. The issue of uncertain prevalence estimate is also applicable to conventional method. Empirical studies on conventional prevalence adjustment also estimated the prevalence from the sample that was subsequently used to assess marker performance - effectively taking its value as known2,3. Further, our explorations were for when performance metrics are derived from a single source population. Without knowing how populations differ from each other, the choice of transportation method will require assumptions on the distribution of causes. On the other hand, when contingency tables from multiple populations are at hand, this choice can be learned from the data. This results in SCC-based model specifications for meta-analysis of marker performance studies, where the estimands are the parameters that govern the joint distribution of $T$, $U$, and $V$. This will provide an alternative to modeling the joint distribution of prevalence, sensitivity, and specificity (or prevalence and predictive values)32,33. The more explanatory nature of SCC-based specification might provide better fit to the data. In addition to pooled estimates of performance metrics and their predictive distribution for a new population, this approach provides an overall estimate of the degree by which universal and alternative causes are responsible for between-population variations, which might be of secondary interest.
Appendix 1: Operationalizing the three types of transportation
Let {TP,FP,FN,TNTP,FP,FN,TN} be the elements of the two-by-two contingency table from the source population. We would like to transport this marker to a new setting where prevalence is $\pi$. Our goal is to construct a predicted contingency table for the target population, defined by $TP^{\prime},FP^{\prime},FN^{\prime},TN^{\prime}$}.
A. Transportation by accuracy:
  1. 1.
    Compute sensitivity and specificity from the source population
    $SE=TP/(TP+FN)$, $SP=TN/(FP+TN)$
  2. 2.
    Combine the SE and SP from the source population with prevalence from the target population
    $TP^{\prime}=\pi SE,\qquad FP^{\prime}=(1-\pi)(1-SP),\qquad FN^{\prime}=\pi(1-SE),\qquad TN^{\prime}=(1-\pi)SP$
B. Transportation by predictive values:
  1. 1.
    Compute positive and negative predictive values from the source population
    $PPV=TP/(TP+FP)$, $NPV=TN/(FN+TN)$
  2. 2.
    Compute $p$, the probability of positive marker results in the new population. Given that $\pi=pPPV+(1-p)(1-NPV)$, we have $p=(\pi-1+NPV)/(PPV+1-NPV)$.
  3. 3.
    Combine the PPV and NPV from the source population with test positivity from the target population
    $TP^{\prime}=pPPV,\qquad FP^{\prime}=p(1-PPV),\qquad FN^{\prime}=(1-p)(1-NPV),\qquad TN^{\prime}=(1-p)NPV$
C. Transportation by the proportional-odds assumption:
  1. 1.
    Map from {TP,FP,FN,TN} to {T,U,V}:
    $T=TP+FP,\qquad U=TP/(TP+FP),\qquad V=FN/((1-T)U)$
  2. 2.
    Find the odds-ratio $x$ such that when applied to $T$, $U$, $V$, results in updated probabilities $T^{\prime}$, $U^{\prime}$, $V^{\prime}$ that correspond to the desired prevalence. Note that prevalence is equal to $T^{\prime}(U^{\prime}+V^{\prime}-U^{\prime}V^{\prime})$.
Of note: $x$ can be found as the real root of the cubic equation $ax^{3}+bx^{2}+cx+d=0$ with
$a$ $=\pi(TUV-TU-TV-UV+T+U+V-1),$
$b$ $=\pi(-3TUV+2TU+2TV+2UV-T-U-V),$
$c$ $=(3\pi-2)TUV-\pi(TU+TV+UV)+TU+UV,$
$d$ $=(1-\pi)TUV;$
which can be solved using standard methods.
  1. 3.
    Recalculate the contingency table from $T^{\prime}$, $U^{\prime}$, $V^{\prime}$:
    $TP^{\prime}=T^{\prime}U^{\prime},\qquad FP^{\prime}=T^{\prime}(1-U^{\prime}),\qquad FN^{\prime}=(1-T^{\prime})U^{\prime}V^{\prime},\qquad TN^{\prime}=1-TP^{\prime}-FP^{\prime}-FN^{\prime}$

### v2

@@
 However, empirical findings indicate that markers often exhibit varying performances across populations.
-For prognostic markers whose results are used to quantify of the risk of an outcome, oftentimes a form of updating is required when the marker is transported to populations with different disease prevalences.
+For prognostic markers that are advertised as predictive risk equations, oftentimes a form of updating is required when the equation is transported to populations with different disease prevalences.
 Here, we revisit transportability of prognostic markers through the lens of the foundational framework of sufficient component causes (SCC).
-We argue that transporting a marker “as is” implicitly assumes predictive values are transportable, whereas conventional prevalence-adjustment shifts the locus of transportability to accuracy metrics (sensitivity and specificity).
-Using a minimalist SCC framework that decomposes risk prediction into its causal constituents, we show that both approaches rely on strong assumptions about the stability of cause distributions across populations.
+We argue that transporting a marker “as is” implicitly assumes predictive values are transportable, whereas conventional prevalence adjustment shifts the locus of transportability to accuracy metrics (sensitivity and specificity).
+Using a minimalist SCC framework that decomposes risk prediction into its causal constituents, we show that both approaches rely on strong assumptions about the stability of cause distributions.
 A SCC framework instead invites making transparent assumptions about how different causes vary across populations, leading to different transportation methods.
 For example, in the absence of any external information other than disease prevalence, a cause-neutral perspective can assume all causes are responsible for change in prevalence, leading to a new form of marker transportation.
-Numerical experiments demonstrate that different transportability assumptions lead to varying degrees of information loss, depending on how population differ from each other in the distribution of causes.
+Numerical experiments demonstrate that different transportability assumptions lead to varying degrees of information loss, depending on the distribution of causes.
 A SCC perspective challenges common assumptions and practices for marker transportability, and proposes transportation algorithms that reflect our knowledge or assumptions about how causes vary across populations.
@@
 Background
-The underlying premise of reporting on the performance of biomarkers, tests, and prediction models (which we generally refer to as “markers”) is that performance metrics are assumed to be transportable from one population to another.
-In this context, we define transportability as conditions under which certain indices of marker performance remain the same in the new population as in the source population (where it was developed).
+The underlying premise of reporting on the performance of biomarkers, tests, and prediction models (which we generally refer to as “markers”) is that performance metrics are transportable from one population to another.
+In this context, we define transportability as conditions under which certain indices of marker performance remain the same across populations.
 Examples of such indices include accuracy (sensitivity and specificity), predictive values, discrimination, and calibration.
@@
 In another example, the performance of machine learning models for detecting pneumonia on chest X-rays substantially declined in data from settings not used to train the model4.
-Often times, when transporting a marker for an outcome to a new population, some information about the characteristics of that population - colloquially referred to as ‘case-mix’- is available.
-One of the most common information about the new population is the outcome prevalence.
+Often times, when transporting a marker for an outcome to a new population, some information about the characteristics of that population - colloquially referred to as case-mix- is available.
+One of the most common information pieces is the outcome prevalence.
 For example, cancer registries often provide a good estimate of cancer risk in a population.
-This poses the challenge as to how such information can be used to revise our assessment of marker performance.
+This poses the challenge as to how such information can be used to revise our assessment of marker performance when transporting it to a new population.
 For risk predictions models that return a quantitative estimate of disease (or outcome) risk given patient characteristics, some form of model updating (or revision) is often needed to correct for the under-estimation or over-estimation of risks in the new population5.
 When the model is a logistic regression, the most basic form of such updating involves modifying the intercept of the model to account for difference in prevalence between the source and target populations1,6,7.
-The underlying justification for this approach is that differences in outcome prevalence are mostly due to factors that act independently of predictors used in the model6.
-Because in a logistic regression model, changing the intercept is equivalent to applying a correcting odds-ratio to predicted risks, this method can be generalized to correcting the outputs of any algorithm that predicts the risk of a binary outcome, including black-box (e.g., machine learning) models8.
+The underlying justification for this approach is that differences in outcome prevalence are mostly due to factors that are distributed, and act, independently of predictors used in the model6.
+Because in a logistic regression model, changing the intercept is equivalent to applying an odds-ratio to predicted risks, this method can be generalized to applying a correcting odds-ratio to the outputs of any algorithm that predicts the risk of a binary outcome, including black-box (e.g., machine learning) models8.
 The transportability of predictive information across populations is an active area of research in predictive analytics and machine learning.
@@
 SCC is a foundational model in that it establishes fully deterministic relationships between causes and effects (rather than a representation of statistical dependencies as in causal graphs).
-In this framework the variability in the outcome is due to the variability of, and interactions between, causes.
 This framework has recently been applied to explore the biologic plausibility of different link functions for modeling binary outcomes, resulting in proposals for more transportable measures of association such as the ‘switch relative risk’16 and more biology-aligned statistical models such as regression by composition17.
-In this paper, we use a simple SCC model to study the most basic prediction setup: a binary factor that is used as a marker for the risk of an outcome.
+In this paper, we use a simple SCC model to study the most basic prediction setup: a binary factor that is used as a marker for the risk of a binary outcome.
 As a reference, we formulate this setup for prognostic markers, where adjustment for outcome prevalence seems to be a topical issue.
-We discuss how logical symmetries can be used to arrive at interpretations applicable to diagnostic markers.
-We create minimalist logic circuits to model a prognostic marker that provides partial, inconclusive information about an outcome of interest.
+We later discuss how symmetries can be used to investigate the transportability of diagnostic markers but leave deeper investigations to future work.
 This setup is the used to study marker transportability across populations, particularly as it relates to the variability in outcome prevalence.
@@
   • The latent binary variable $V$ represents all alternative causes - for example pathways related to the effect of tobacco smoking, which increases the risk of breast cancer even among those without a BRCA1 mutation.
-The absence of $V$ is responsible for false-negative marker values.
+The presence of $V$ is responsible for false-negative marker values.
 We assign the values of 1 and 0, respectively, to the ‘on’ and ‘off’ status of each of these switches.
 For the brevity of notations, by a simple character, we refer to the ‘on’ value, and by its dot-accented to its negated value (e.g., $\dot{U}=\neg U$ indicating $U=0$).
-We use the shorthand notation $P(X)$ for $P(X=1)$.
-For example, $P(T\dot{V})=P(T\dot{V}=1)=P(T=1,V=0)$.
 The above setup results in the following outcome-generating process for prognostic markers:
-$D=(T\land U)\lor(V\land U)=(T\lor V)\land U.$
+$D=(T\land U)\lor(V\land U)=(T\lor V)\land U,$
+where $\land$ and $\lor$ are logical AND and OR, respectively.
+Given that $T$, $U$, and $V$ are all binary, a population is made up of 8 subgroups.
+As $U$ and $V$ are latent variables, the observed properties of the marker in a population is manifested in terms of $P(T,D)$, i.e., the two-by-two (contingency) table of marker by outcome status probabilities.
+We represent the contingency table by the sequence TP,FP,FN,TN{TP,FP,FN,TN}, where $TP=P(TD=1),FP=P(T\dot{D}=1),FN=P(\dot{T}D=1),TN=P(\dot{T}\dot{D}=1)$ are, respectively, true positive, false positive, false negative, and true negative probabilities.
+Our interest is in the four fundamental performance characteristics of binary markers: positive predictive value (PPV=P(D=1|T=1)=TP/(TP+FP){PPV}=P(D=1|T=1)=TP/(TP+FP)), negative predictive value (NPV=P(T=1|D=0)=TN/(FN+TN){NPV}=P(T=1|D=0)=TN/(FN+TN)), sensitivity (SE=P(T=1|D=1)=TP/(TP+FN){SE}=P(T=1|D=1)=TP/(TP+FN)), and specificity (SP=P(T=0|D=0)=TN/(FP+TN){SP}=P(T=0|D=0)=TN/(FP+TN)), as a function of prevalence ($P(D=1)=TP+FN$).
+By convention PPV and NPV are referred to as predictive values and SE and SP as accuracy metrics.
 Before proceeding, we note that one can create an equally parsimonious SCC setup by swapping the logical AND and OR in the above equation, resulting in the setup $D=(T\land V)\lor U$.
@@
 This becomes obvious by considering that the complementary marker (whose positive and negative results are swapped) is a marker for not experiencing the outcome under the reference setup: $\dot{D}=(\dot{T}\lor\dot{V})\land\dot{U}$.
-Because of this, any pattern we observe for PPV is also observed for NPV in this alternative formulation, albeit in the opposite direction, and vice versa.
+Because of this, any pattern we observe for PPV in the main formulation is also observed for NPV in this alternative formulation, albeit in the opposite direction, and vice versa.
 Similarly, any pattern we observe for SE in the reference setup is also observable, in the opposite direction, for SP in this alternative setup.
@@
 Top: logic equation; Middle: SCC diagram; Bottom: The resulting contingency table
-Reference setup Reciprocal setup
+Reference setup Symmetrical setup
 $D=(T\lor V)\land U$ $D=(T\land V)\lor U$
@@
 [Uncaptioned image] [Uncaptioned image]
-Given that $T$, $U$, and $V$ are all binary, a population is made up of 8 subgroups.
-For example, $T\dot{U}V$ is the subgroup where $T=1$, $U=0$, and $V=1$.
-As $U$ and $V$ are latent variables, the observed properties of the marker in a population is manifested in terms of $P(T,D)$, i.e., the two-by-two (contingency) table of marker by outcome status probabilities in that population.We express a contingency table by the sequence TP,FP,FN,TN{TP,FP,FN,TN}, where $TP=P(TD),FP=P(T\dot{D}),FN=P(\dot{T}D),TN=P(\dot{T}\dot{D})$.
-If all we know about populations is their contingency tables, the full joint distribution $P(T,U,V)$ is unidentifiable, given its specification requires seven degrees of freedom but the contingency table only provides three.
-However, any joint distribution can be replaced by independent Bernoulli distributions for $P(T)$, $P(U)$, and $P(V)$ that would result in the same contingency table.
-Given specifying three Bernoulli distributions requires three degrees of freedom, this setup is identifiable and has a unique (1:1) mapping to a given contingency table.
-Because our focus is on metrics that are determined by the contingency table, without loss of generality the following we assume independence i.e., $P(T,U,V)=P(T)P(U)P(V)$.
-Our interest is in the four fundamental performance characteristics of binary markers: sensitivity (SE=P(T|D)=TP/(TP+FN){SE}=P(T|D)=TP/(TP+FN)), specificity (${SP}=P(\dot{T}|\dot{D})=TN/(FP+TN)$), positive predictive value (PPV=P(D|T)=TP/(TP+FP){PPV}=P(D|T)=TP/(TP+FP)), and negative predictive value (${NPV}=P(\dot{T}|\dot{D})=TN/(FN+TN)$), as a function of prevalence ($P(D)=TP+FN$).
-By convention by SE and SP we refer to accuracy metrics, and by PPV and NPV as predictive values.
-As a numerical example, consider a population where $P(T)=$ 0.25, $P(U)=$ 0.75, and $P(V)=$ 0.50.
-The 8 subgroups are scattered into the four cells of the contingency table (see Table 1).
-The contingency table emerging from this setup is {0.188,0.062,0.281,0.469}.This results in a prevalence ($P(D)$) of 0.469, SE of 0.400, SP of 0.882, PPV of 0.750, and NPV of 0.625.
+  • Subgroups are defined as the combination of causes such that the product evaluates to 1. For example, $T\dot{U}V$ is the subgroup where $T=1$, $U=0$, and $V=1$.
+Mapping between contingency table and cause probabilities
+If all we know about marker performance in a populations is its contingency table, the full joint distribution $P(T,U,V)$ is unidentifiable, given its specification requires seven degrees of freedom but the contingency table only provides three.
+However, any joint distribution can be replaced by independent marginal distributions for $P(T=1)$, $P(U=1)$, and $P(V=1)$ that would result in the same contingency table.
+We use the shorthand notations $P_{T}=P(T=1),P_{U}=P(U=1)$, and $P_{V}=P(V=1)$.
+Given specifying three Bernoulli distributions for these probabilities requires three degrees of freedom, this setup is identifiable and has a unique (1:1) mapping to a given contingency table.
+This 1:1 mapping can thus be formulated as follows:
+  • From {TP,FP,FN,TNTP,FP,FN,TN} to $P_{T},P_{U},P_{V}$}:
+    $P_{T}=TP+FP,\quad P_{U}=TP/(TP+FP),\quad P_{V}=FN/[(1-P_{T})P_{U}]$
+  • From $P_{T},P_{U},P_{V}$} to {TP,FP,FN,TNTP,FP,FN,TN}:
+    $TP=P_{T}P_{U},\quad FP=P_{T}(1-P_{U}),\quad FN=(1-P_{T})P_{U}P_{V},\quad TN=1-TP-FP-FN.$
+As a numerical example, consider a population where $P_{T}=$ 0.25, $P_{U}=$ 0.75, and $P_{V}=$ 0.50.
+The 8 subgroups are scattered into the four cells of the contingency table (see Table 1), which is presented in the top row of Table 2.
+This results in a prevalence of 0.469, SE of 0.400, SP of 0.882, PPV of 0.750, and NPV of 0.625.
 We will use this simple numerical example as a case study on various ways a marker can be transported.
+Table 2: Numerical example of different marker transportation methodology
+Description { $P_{T},P_{U},P_{V}$ } Contingency table SE SP PPV NPV
+Source population {0.250,0.750,0.500} {0.188,0.062,0.281,0.469} 0.400 0.882 0.750 0.625
+Target population {0.333,0.800,0.667} {0.267,0.067,0.356,0.311} 0.429 0.824 0.800 0.467
+By predictive values {0.333,0.750,0.500} {0.250,0.083,0.250,0.417} 0.500 0.833 0.750 0.625
+By accuracy {0.293,0.848,0.623} {0.249,0.044,0.373,0.333} 0.400 0.882 0.848 0.472
+Proportional odds {0.349,0.829,0.617} {0.290,0.060,0.333,0.318} 0.465 0.841 0.829 0.489
+SE: sensitivity; SP: specificity; PPV: positive predictive value; NPV: negative predictive value
+  • Contingency table presents the sequence TP (true positive), FP (false positive), FN (false negative), TN (true negative)
 Two common methods of transportability
 In this section we review the two conventional approaches for transporting a marker to a new population with potentially different outcome prevalence: transportation via accuracy (SE & SP), and transportation via predictive values (PPV & NPV).
-In transportation by accuracy, we assume the SE and SP are equal between the source and target populations.
-The contingency table in the target population can be recovered by knowing outcome prevalence in the target population, denoted by $\pi$.
-For example, the proportion of true positives consists of those with the disease who also rest positive: $TP=\pi\times SE$, and so on.
-From this contingency table, various performance metrics can be derived.
-Details of this approach are provided in Appendix 1 (section A).
+Ultimately, using a marker for decision-making entails interpreting its predictive values - i.e., $P(D|T)$.
+Different transportation methods construct these predictive values by combining different pieces of information from the source and target populations.
+Transportation by predicted values
+Transportation by predictive values refers to the where predictive values from the source population are considered to be the same as in the target population.
+Given that $P(T,D)=P(T)P(D|T)$, this transportability means our estimate of the contingency table in the target population is determined by $P(T)$ in that population, combined with predictive values from the source population.
+This approach is more common for prognostic markers, a typical example being risk scoring tools that directly return an estimate of $P(D=1|T)$ for someone with marker value $x$.
+For a binary marker, this risk equation can be written as P(D=1|T)=(1NPV)+[PPV(1NPV)]TP(D=1|T)=(1-{NPV})+[{PPV}-(1-{NPV})]T.
+In our numerical example, advertising prognostic information as a risk equation would result in $P(D=1|T)=0.375+0.375T$.
+Transportation by accuracy
+In transportation by accuracy, we assume that $P(T|D)$, i.e., the SE and SP, are equal between the source and target populations.
+In order to construct predictive probabilities, we apply the Bayes’ rule:
+$P(D=1|T)=\frac{P(D=1)P(T|D=1)}{P(T)}=\frac{P(D=1)P(T|D=1)}{P(D=1)P(T|D=1)+P(D=0)P(T|D=0)}.$
+Thus, under this transportability scheme we need access to prevalence in the target population to derive predictive values.
 This approach is the modus operandi for binary diagnostic markers, where the Bayes’ rule is used to combine pre-test probability with marker accuracy estimates to derive the post-test probability of the outcome21,22.
 Without any information that would distinguish the individual under evaluation, the pre-test probability is taken to be disease prevalence in the target population, which is equal to this mode of transportation.
-Transportation by predictive values refers to the where we assume it is the PPV and NPV from the source population that are transportable.
-Under this assumption, differences in outcome prevalence is due to the change in the prevalence of marker-positive individuals across populations.
-Again, if the target prevalence is known, PPV and NPV can be used to recover the contingency table in the target population.
-Details of this approach are provided in Appendix 1 (section B).
-This approach is more common for prognostic markers, a typical example being risk scoring tools that directly return an estimate of $P(D|T=x)$ for someone with marker value $x$.
-For a binary marker, this risk equation can be written as P(D|T=x)=(1NPV)+[PPV(1NPV)]xP(D|T=x)=(1-{NPV})+[{PPV}-(1-{NPV})]x.
-This equation indicates the implied assumption that the predictive values of the marker are transportable.
-In our numerical example, advertising prognostic information as a risk equation would result in $P(D|T)=0.375+0.375T$.
-Returning to our numerical example, consider a population with 25% higher prevalence of breast cancer ($P(D)=0.586$).
-Transportation by accuracy results in a contingency table of {0.234,0.049,0.352,0.365}, corresponding to a PPV of 0.828 and NPV of 0.510.
-On the other hand, transportation by predictive values gives rise to a contingency table of {0.422,0.141,0.164,0.273}, corresponding to a SE of 0.720 and SP of 0.660.
-Prevalance adjustment
-A common way that prognostic markers are transported across populations is in the format of a risk equation for $P(D|x)$, either explicitly, as is the case in regression-based prediction models, or implicitly, as in black-box (e.g., machine learning) models.
-Consider applying the Bayes’ rule to the outcome risk conditional on marker value:
-$P(D|T=x)=\frac{P(D)P(T=x|D)}{P(x)}=\frac{P(D)P(T=x|D)}{P(D)P(x|D)+P(\dot{D})P(x|\dot{D})}.$
+Returning to our running example, imagine we are transporting the above-mentioned prognostic marker to a population where $P_{T}=$ 0.250, $P_{U}=$ 0.750, $P_{V}=$ 0.500.
+The true contingency table and marker performance metrics are provided in Table 2.
+Transportation by predictive values preserves PPV and NPV in the target population, but will result in SE and SP of 0.5 and 0.833.
+Now, imagine we know the target population prevalence, which is 0.622 (32.7% higher than in the source population).
+If we consider transportation by accuracy, we will arrive at the PPV and NPV values of, respectively, 0.750 and 0.625.
+Prevalence adjustment for risk equations
+A common way that prognostic markers are transported across populations is in the format of a risk equation for $P(D|T)$, either explicitly, as is the case in regression-based prediction models, or implicitly, as in black-box (e.g., machine learning) models.
 A familiar modeling framework for binary outcomes is the logistic regression.
 There, the logit function ($\mbox{logit}(x)=\log(\frac{x}{1-x})$) is used as link function connecting marker value to outcome probability.
-Applying this function to both sides of the risk equation for marker values, we have
-$\mbox{logit}(P(D|T=x))=\mbox{logit}(P(D))+\log\left(\frac{P(T=x|D=1)}{P(T=x|D=0)}\right).$
+Applying this function to both sides of the previous equation, we have
+$\mbox{logit}(P(D=1|T))=\mbox{logit}(P(D=1))+\log\left(\frac{P(T|D=1)}{P(T|D=0)}\right).$
 The last term on the right-hand side is the likelihood ratio (LR) of the marker between the diseased and non-diseased groups.
@@
 For the marker to be transportable, both these LRs need to remain constant across populations, which will be the case if and only if both SE and SP remain constant.
-This in turn implies that prevalence adjustment for binary markers is equal to considering SE and SP to be transportable.
-Under what conidtions are performance metrics transportable?
-The above reasoning shows why adjustment for prevalence is generally expected to improve the transportability of a marker, as the conventional wisdom is that SE and SP, being defined within the diseased and non-diseased groups, are less dependent on prevalence23.
+This in turn implies that prevalence adjustment for binary markers is equal to transportation by accuracy; i.e., considering SE and SP to be transportable.
+Under what conditions are performance metrics transportable?
+The above reasoning shows why adjustment for prevalence is generally expected to improve transportability, as the conventional wisdom is that SE and SP, being defined within the diseased and non-diseased groups, are less dependent on prevalence23.
 While this might be intuitive as a general observation, under the SCC framework, none of these metrics are truly intrinsic.
@@
 Metric Transportability condition
-SE{SE} ${P(\dot{T}UV)}/{P(TU)}=c$
-SP{SP} ${P(T\dot{U})}/({P(\dot{T})-P(\dot{T}UV)})=c$
-PPV{PPV} $P(U\mid T)=c$
-NPV{NPV} $P(UV\mid\dot{T})=c$
+PPV{PPV} $P(U=1|T=1)=c$
+NPV{NPV} $P(UV=1|T=0)=c$
+SE{SE} ${P(T=1|U=1)}/{P(\dot{T}V=1|U=1)}=c$
+SP{SP} ${P(T\dot{U}=1)}/[{P(T=0)-P(\dot{T}UV=1)}]=c$
 SE: sensitivity; SP: specificity; PPV: positive predictive value; NPV: negative predictive value
-  • As an example of derivations, consider PPV. Its transportability means $P(D|T)=c$ (a constant). But $P(D|T)=(P(TU\dot{V})+P(TUV))/P(T)=P(TU)/P(T)=P(U|T)$.
+  • As an example of derivations, consider PPV. Its transportability means $P(D=1|T=1)=c$ (a constant). But $P(D=1|T=1)=(P(TU\dot{V}=1)+P(TUV=1))/P(T=1)=P(TU=1)/P(T=1)=P(U=1|T=1)$.
 For the PPV to be transportable, the prevalence of universal causes among marker-positive individuals should remain constant.
@@
 In our breast cancer example, this would mean populations vary in breast cancer prevalence only because they differ in the BRCA1 mutation rate.
-For many markers, including BRCA1, this assumption is unrealistic, as many factors contribute to the variability in breast cancer risk24.
+For many outcomes, including breast cancer, this assumption is unrealistic, as many factors contribute to the variability in breast cancer risk24.
 In comparison, conditions for the transportation by accuracy are more complicated.
@@
 SE: sensitivity; SP: specificity; PPV: positive predictove value; NPV: negative predictive value
-  • The base setup is $P(T)=0.25$, $P(U)=0.75$, $P(V)=0.5$. Changes across populations are modeled by applying varying odds-ratios to the components that vary (labeled on top of each panel).
+  • The base setup is $P_{T}=0.25$, $P_{U}=0.75$, $P_{V}=0.5$. Changes across populations are modeled by applying varying odds-ratios to the components that vary (labeled on top of each panel).
 Other forms of marker transportation under the SCC framework
@@
 Consider investigating a rare ancestral mutation as a prognostic marker for cancer that is known to have minimal variations across populations.
-In this case, it is reasonable to assume that $P(T)$ remains essentially constant across populations.
+In this case, it is reasonable to assume that $P_{T}$ remains essentially constant across populations.
 When transporting this marker, our focus would therefore be on scenarios where differences in $P(U,V)$ derive variations in cancer prevalence.
@@
 Indeed, both these transportation methods implicitly require that an increase in outcome prevalence be accompanied by change in marker positivity.
+Proportional-odds transportability method
 But, what if the only information we have is differences in outcome prevalence across populations?
@@
 For transportability by predictive values, one can question as to why only one cause ($T$, the one we happen to be measuring) should be responsible for change in prevalence.
-Transportability by SE and SP is also questionable due to unclear assumptions it places on the distribution of causes.
-For this general case, it might be more reasonable to take an explicit yet neutral stance about the variations in causes that bring about variations in prevalence, for example that all causes move by the same extent to cause change in prevalence - this approach makes no exception for thae causes for the mere fact that we are measuring it.
-As an implementation of this approach, one can solve for a common odds-ratio that, when applied to $P(T)$, $P(U)$, and $P(V)$ in the source population, results in a new population that matches the outcome prevalence in the target population.
+Transportability by accuracy is also questionable due to unclear assumptions it places on the distribution of causes.
+For this general case, it might be more reasonable to take a neutral yet transparent stance about the variations in causes, for example that all causes move by the same extent to cause change in prevalence.
+As an implementation of this approach, instead of applying a correcting odds-ratio to predicted values, one can solve for a common odds-ratio that, when applied to $P_{T}$, $P_{U}$, and $P_{V}$ in the source population, results in a new population that matches the outcome prevalence in the target population.
 This “proportional odds” assumption is a new, distinct transportability approach that would generate different updated values of marker performance metrics compared with both conventional methods.
-Details of operationalizing this approach are provided in Appendix 1 (section C).
-In our numerical example, transporting the marker to a new population with 25% higher prevalence would require applying an odds-ratio of 1.434 to the three causes.
-This odd-ratio can be derived via standard univariate root finding (given that there is a monotonical relationship between this odds-ratio and prevalence).
-It can also be expressed as the real root of a cubic equation (see Appendix 1 - section C).
-The contingency table under this method of transportability will be {0.262,0.061,0.323,0.353}.
-This in turn results in SE of 0.448 and SP of 0.853.
-The PPV and NPV are, respectively, 0.811 and 0.522.
-Formulating marker performance as a risk equation would thus result in $P(D|T)=0.189+0.333T$).
+Details of operationalizing this approach are provided in Appendix 2.
+In our numerical example (last row of Table 2), we need to apply an odds-ratio of 1.612 to the three causes to match the target population’s outcome prevalence.
+This odd-ratio can be derived via standard univariate root finding - given that there is a monotonical relationship between this odds-ratio and prevalence.
+It can also be expressed as the real root of a cubic equation for which standard solutions are available (see Appendix 2).
+The contingency table under this method of transportability will be {0.290, 0.060, 0.333, 0.318}.
+This in turn results in SE of 0.465 and SP of 0.841.
+The PPV and NPV are, respectively, 0.829 and 0.489.
+Formulating marker performance as a risk equation would thus result in $P(D|T)=0.171+0.317x$.
 Note that both the intercept and slope (coefficient for $T$) of the risk equation are updated, which is different from the risk equation under the assumption of transportability of predictive values (where both remain unchanged), or accuracy (where only the intercept is updated).
@@
 The discrepancy between an assumed versus the true marker performance can be measured in different ways, each focusing on different consequences of such discrepancy (for example, through measures of discrimination, calibration, or prediction error).
-A more foundational approach is to measure information loss according to information theory, which in this context can be measured via the Kullback-Leibler divergence ($D_{KL}$)25.
+A more foundational approach is to measure information loss according to information theory.
+In this context, the relevant measure would be the Kullback-Leibler divergence ($D_{KL}$)25.
 This information-theoretic measure quantifies the discrepancy between a true distribution and a candidate distribution in terms of ‘excess surprise’, which can be interpreted as the additional number of bits required to encode information from the true distribution using the candidate distribution, rather than using the true distribution itself26.
@@
 We modeled population-generating scenarios similar to those presented in Figure 1.
-Because in realistic settings all three causes are likely to vary, we explored this setup with more depth.
+Because in many realistic settings all three causes are likely to vary, we explored this setup with more depth.
 We modeled scenarios where changes in the three causes had different strengths of correlation with each other (ranging from changes being uncorrelated to fully correlated).
-We also modeled a ‘maximum entropy’ scenario where the prevalence of the eight subgroups follows Dirichlet(1,1,1,1,1,1,1,1); this distribution has maximum information-theoretic randomness (entropy)27.
+We also modeled a ‘maximum entropy’ scenario where the three causes have independent uniform distributions; albeit unrealistic, this scenario represents maximum theoretical randomness across populations (therefore, marker performance is entirely untransportable)27.
 For each scenario, we simulated random pairs of source and target populations according to a given population-generating mechanism.
+When a cause was made variable, it was assumed that the logit-transformed probability of that cause had a standard Normal distribution (e.g., $\mbox{logit}(P_{T})\sim\mbox{Normal}(0,1)$ when where $P_{T}$ was variable across populations).
 For each pair, we applied the following three transportation strategies: by accuracy (SE and SP), by predictive values (PPV and NPV), and via the proportional-odds assumption explained previously.
@@
 This confirms that if the outcome prevalence variation is entirely due to $T$, PPV and NPV are fully transportable ($D_{KL}=0$).
-However, under other scenarios, transporting by predictive values resulted in substantial information loss.
-Transporting by accuracy reduced information loss in some scenarios, albeit in none was the loss zero, indicating that none of the modeled scenarios were compatible with stable SE and SP.
-Transporting by accuracy performed worse than transporting by predictive values in some scenarios, indicating that prevalence adjustment is not guaranteed to improve the transportability of a prognostic marker.
-The proportional-odds assumption was also reduced information loss in some scenarios, but increased it in others.
-For the scenarios where the three causes changed together, transportation by accuracy generally reduced information loss compared with transportation by predictive values (unless the correlation among causes were weak).
-In all those situations, transportation via he proportional-odds assumption performed better than the conventional transportability methods.
-As expected, in the scenarios where changes in causes were positively correlated.
+However, under other scenarios, transportation by predictive values resulted in information loss, sometimes substantially.
+Transportation by accuracy reduced information loss in some scenarios, but increased in others.
+This indicates that prevalence adjustment is not guaranteed to improve the transportability of a prognostic marker.
+In none of the scenarios was the loss zero, indicating that none of the modeled scenarios were compatible with stable SE and SP - again, due to the complex requirements for the stability of these metrics.
+The proportional-odds assumption also reduced information loss in some scenarios, but increased it in others.
+For the scenarios where the three causes changed together, transportation by accuracy generally reduced information loss compared with transportation by predictive values.
+In such scenarios, transportation via the proportional-odds assumption performed better than both conventional transportability methods.
 Figure 2: Information loss associated with different methods for transporting a marker under various population-generating scenarios
@@
   • Kullback-Leibler divergence $(D_{KL})=\sum_{i=1}^{4}P(a_{i})\log_{2}(P(b_{i})/P(a_{i}))$ with $a_{1},a_{2},a_{3},a_{4}$ being the four cells of the contingency table from the correct $P(T,D)$, and $b_{i}$s the corresponding ones from $P(T,D)$ constructed given the transportability method.
-  • The base population is as follows: $P(T)=0.25$, $P(U)=0.75$, and $P(V)=0.5$.
-When a cause varies, it is given a logit-normal distribution with mean being equal to the base population’s point estimate, and standard deviaioon of 0.1.
-  • No coorelation, weak, moderate, strong, and full correlation correspond to correlation coefficients of, respectively, 0, 0.25, 0.50, 0.75, amd 1 for logit-transformed probabilities.
+  • Aside in the maximum entropy scenario (whose description provided in text), he variable component in each scenario had a standard logit-normal distribution e.g., $\mbox{logit}(T)\propto\mbox{Normal}(0,1)$ for scenarios where $T$ was variable.
+  • No correlation, weak, moderate, strong, and full correlation correspond to correlation coefficients of, respectively, 0, 0.25, 0.50, 0.75, amd 1 for logit-transformed probabilities.
 Discussion
@@
 However, in scenarios where causes were positively correlated, which is likely to be common, transporting by accuracy did reduce information loss compared with transportability by predictive values.
-These results are based on a selected set of population-generating mechanisms and a single set of core parameters for the SCC model ($P(T)=0.25$, $P(U)=0.75$, $P(V)=0.5$).
+However, in the same scenarios, the SSC-based ‘proportional odds’ method of transportation performed better.
+Our simulations were related to a selected scenarios and particular ways the causes vary across populations.
 Still, they should be sufficient to question some common assumptions and practices.
@@
 While these insights are from studying binary markers, the core findings can be extended to continuous markers and multi-variable risk equations.
-Transforming such markers ‘as is’ is equal to assuming that stratum-specific predictive values ($P(D|T=x)$) remain constant.
-This is equal to attributing variations in prevalence entirely to variations in $P(T)$ (aka case-mix).
+Transforming such markers ‘as is’ is equal to assuming that stratum-specific predictive values ($P(D|T)$) remain constant.
+This is equal to attributing variations in prevalence entirely to variations in $P_{T}$ (aka case-mix).
 Consider a multi-variable risk score such as the QRISK for cardiovascular diseases28.
-Even though this model includes up to 22 predictors, is it plausible that other causal pathways to cardiovascular diseases that might vary across populations, including yet-to-be-discovered genetic risk factors, environmental exposures, subtle lifestyle choices, access to and quality of preventive care, and so on, are the same across populations?
+Even though this model includes up to 22 predictors, is it plausible that other causal pathways to cardiovascular diseases that might vary across populations, including environmental exposures, subtle lifestyle choices, access to and quality of preventive care, and yet-to-be-discovered genetic risk factors, are the same across populations?
 This assumption is needed for claiming that QRISK predictions, developed using primary care UK data, are transportable to other settings.
-On the other hand, conventional prevalence adjustment would indicate that stratum-specific likelihood ratio, i.e., $P(x|D)/P(x|\dot{D})$ would be transportable.
-This assumption will hold if the distribution of predicted risks among those who will experience the outcome and those who will not remains is same across diverse populations.
+On the other hand, conventional prevalence adjustment would indicate that stratum-specific likelihood ratio, i.e., $P(x|D=1)/P(x|D=0)$ would be transportable.
+This assumption will hold if the distribution of predicted risks among those who will experience the outcome and those who will not remains the same across diverse populations.
 This condition places very specific constraints on how the causes of diseases should vary across populations.
-We focused on prognostic markers given the recent debates on transportabiltiy of prognostic information and the merit of prevalence-adjustment to improve transportability.
+We focused on prognostic markers given the recent debates on transportability of prognostic information and the merit of prevalence-adjustment to improve transportability.
 The application of SCC framework for diagnostic markers is also important and deserves its own airing.
@@
 This assumption is similar to the assumption of transportability by predictive values for prognostic markers.
-This strong requirement can explain why these metrics tend to vary across populations29–31.
-Further details of how the SCC framework can inform transportability of diagnostic markers needs to be pursued in future work.
+This strong requirement can explain why sensitivity and specificity tend to vary across populations29–31.
+How the SCC framework can inform transportability of diagnostic markers needs to be pursued in future work.
 How can these findings inform practice?
 If reliable, context-specific information is available on the distribution of specific predictors and their relationships with the outcome and with each other, methods based on causal graphs can be used to build transportable models or to design tailored transportation strategies11.
-In contrast, when only general information, such as outcome prevalence or test positivity rates are available, the SCC framework can be used to formulate tailored transportability algorithms that utilize such information.
+In contrast, when only general information, such as outcome prevalence or test positivity rates, are available, the SCC framework can be used to formulate transportability algorithms that utilize such broad information.
 We question the contemporary practice of advertising risk prediction models for outcome risk as transportable.
-We believe risk equations should explicitly mention the outcome risk in their source population, and prevalence adjustment should be advocated as a necessary step before using the equation in a population with different outcome prevalence.
-Yet, while conventional prevalence adjustment may serve as a reasonable default, it cannot be assumed to consistently improve performance.
-Instead of fixed transportability rules applied to common performance metrics, the SCC framework offers a foundation for updating marker performance in ways that are more grounded in mechanistic interactions among potential causes.
+We believe risk equations should explicitly mention the outcome risk in their source population, and adjustment for prevalence should be advocated as a necessary step before using the equation in a population with different outcome prevalence.
+Yet, while conventional prevalence adjustment may serve as a reasonable choice, it cannot be assumed to consistently improve performance.
+Instead of fixed transportability rules, the SCC framework offers a foundation for updating marker performance in ways that are more grounded in mechanistic interactions among potential causes.
 This results in new transportability algorithms based on assumptions on how underlying causes vary, rather than keeping performance metrics, which are emergent properties of such causes, constant.
@@
 It is motivated by the notion that oftentimes differences in outcome prevalence are multi-factorial.
-An agnostic stance that that treats all causes as equally responsible is more defendable than blaming only one (as in transportation by predictive values) or assuming they move in unxplainable ways just to keep sensitivity and specificity constant (as in transportability by accuracy).
+An agnostic stance that treats all causes as equally responsible is more defendable than blaming only one (as in transportation by predictive values) or assuming they move in unexplainable ways just to keep sensitivity and specificity constant (as in transportability by accuracy).
+Interestingly, in our numerical experiments, whenever conventional prevalence-adjustment performed better than transportability by predictive values, the SCC-based proportional-odds method performed better than both, indicating it might be a better choice.
 Of course, variations in this method are also possible.
-Imagine, for example, that in addition to knowing target prevalence, we have knowledge of marker positivity rate ($P(T)$) in the new population (e.g., from an independent cross-sectional study).
-One can then directly adjust $P(T)$ to match this value, and then apply a common odds-ratio to other causes to match target prevalence.
+Imagine, for example, that in addition to knowing target prevalence, we have knowledge of marker positivity rate ($P_{T}$) in the new population (e.g., from an independent cross-sectional study).
+One can then directly adjust $P_{T}$ to match this value, and then apply a common odds-ratio to other causes to match target prevalence.
 We conclude this paper by providing a few areas for further inquiry.
@@
 In addition to pooled estimates of performance metrics and their predictive distribution for a new population, this approach provides an overall estimate of the degree by which universal and alternative causes are responsible for between-population variations, which might be of secondary interest.
-Appendix 1: Operationalizing the three types of transportation
-Let {TP,FP,FN,TNTP,FP,FN,TN} be the elements of the two-by-two contingency table from the source population.
-We would like to transport this marker to a new setting where prevalence is $\pi$.
-Our goal is to construct a predicted contingency table for the target population, defined by $TP^{\prime},FP^{\prime},FN^{\prime},TN^{\prime}$}.
-A. Transportation by accuracy:
+Appendix: Transportation by the proportional-odds assumption
+Let {TP,FP,FN,TNTP,FP,FN,TN} be the elements of the two-by-two contingency table.
+We would like to transport this marker to a new setting where prevalence is $\pi^{*}$.
+Our goal is to construct a predicted contingency table for the target population, defined by $TP^{*},FP^{*},FN^{*},TN^{*}$}.
   1. 1.
-    Compute sensitivity and specificity from the source population
-    $SE=TP/(TP+FN)$, $SP=TN/(FP+TN)$
+    Map from {TP,FP,FN,TNTP,FP,FN,TN} to $P_{T},P_{U},P_{V}$} (see text)
   2. 2.
-    Combine the SE and SP from the source population with prevalence from the target population
-    $TP^{\prime}=\pi SE,\qquad FP^{\prime}=(1-\pi)(1-SP),\qquad FN^{\prime}=\pi(1-SE),\qquad TN^{\prime}=(1-\pi)SP$
-B. Transportation by predictive values:
-  1. 1.
-    Compute positive and negative predictive values from the source population
-    $PPV=TP/(TP+FP)$, $NPV=TN/(FN+TN)$
-  2. 2.
-    Compute $p$, the probability of positive marker results in the new population. Given that $\pi=pPPV+(1-p)(1-NPV)$, we have $p=(\pi-1+NPV)/(PPV+1-NPV)$.
-  3. 3.
-    Combine the PPV and NPV from the source population with test positivity from the target population
-    $TP^{\prime}=pPPV,\qquad FP^{\prime}=p(1-PPV),\qquad FN^{\prime}=(1-p)(1-NPV),\qquad TN^{\prime}=(1-p)NPV$
-C. Transportation by the proportional-odds assumption:
-  1. 1.
-    Map from {TP,FP,FN,TN} to {T,U,V}:
-    $T=TP+FP,\qquad U=TP/(TP+FP),\qquad V=FN/((1-T)U)$
-  2. 2.
-    Find the odds-ratio $x$ such that when applied to $T$, $U$, $V$, results in updated probabilities $T^{\prime}$, $U^{\prime}$, $V^{\prime}$ that correspond to the desired prevalence.
-Note that prevalence is equal to $T^{\prime}(U^{\prime}+V^{\prime}-U^{\prime}V^{\prime})$.
+    Find the odds-ratio $x$ such that when applied to $P_{T}$, $P_{U}$, $P_{V}$, results in updated probabilities $P_{T}^{*}$, $P_{U}^{*}$, $P_{V}^{*}$ that correspond to the desired prevalence.
+Note that prevalence is equal to $P_{U}^{*}(P_{T}^{*}+P_{V}^{*}-P_{T}^{*}P_{V}^{*})$.
 Of note: $x$ can be found as the real root of the cubic equation $ax^{3}+bx^{2}+cx+d=0$ with
-$a$ $=\pi(TUV-TU-TV-UV+T+U+V-1),$
-$b$ $=\pi(-3TUV+2TU+2TV+2UV-T-U-V),$
-$c$ $=(3\pi-2)TUV-\pi(TU+TV+UV)+TU+UV,$
-$d$ $=(1-\pi)TUV;$
+$a$ $=(1-\pi^{*})P_{T}P_{U}P_{V},$
+$b$ $=(3\pi^{*}-2)P_{T}P_{U}P_{V}-\pi^{*}[P_{T}P_{U}+P_{T}P_{V}+P_{U}P_{V}]+P_{T}P_{U}+P_{U}P_{V},$
+$c$ $=-\pi^{*}[3P_{T}P_{U}P_{V}-2P_{T}P_{U}-2P_{T}P_{V}-2P_{U}P_{V}+P_{T}+P_{U}+P_{V}],$
+$d$ $=-\pi^{*}(1-P_{T})(1-P_{U})(1-P_{V});$
 which can be solved using standard methods.
   1. 3.
-    Recalculate the contingency table from $T^{\prime}$, $U^{\prime}$, $V^{\prime}$:
-    $TP^{\prime}=T^{\prime}U^{\prime},\qquad FP^{\prime}=T^{\prime}(1-U^{\prime}),\qquad FN^{\prime}=(1-T^{\prime})U^{\prime}V^{\prime},\qquad TN^{\prime}=1-TP^{\prime}-FP^{\prime}-FN^{\prime}$
+    Map the resulting $P_{T}^{*}$, $P_{U}^{*}$, $P_{V}^{*}$} to $TP^{*},FP^{*},FN^{*},TN^{*}$} (see text)

### v3

@@
 Abstract
-Transportability, the ability to maintain performance across populations, is a desirable property of of markers of clinical outcomes.
+Transportability, the ability to maintain performance across populations, is a desirable property of markers of clinical outcomes.
 However, empirical findings indicate that markers often exhibit varying performances across populations.
@@
 A SCC perspective challenges common assumptions and practices for marker transportability, and proposes transportation algorithms that reflect our knowledge or assumptions about how causes vary across populations.
-footnotetext: From Faculty of Pharmaceutical Sciences and Faculty of Medicine (MS), Vancouver, Canada; School of Population Health, WHO Collaborating Centre for Climate Change and Health Impact Assessment, and enAble Institute, Faculty of Health Sciences, Curtin University, Perth, Australia (GP); Faculty of Medicine, Universitas Negeri Malang, Indonesia (GP); Saw Swee Hock School of Public Health, National University of Singapore, Singapore (WC)footnotetext: * Correspondence to Mohsen Sadatsafavi, Room 4110, Faculty of Pharmaceutical Sciences, 2405 Wesbrook Mall, Vancouver, British Columbia, V6T1Z3, Canada; email: mohsen.sadatsafavi@ubc.ca
+Keywords: Prognosis; Predictive Value of Tests; Sensitivity and Specificity; Biomarkers; Etiology; Epidemiologic Methods
+footnotetext: From Faculty of Pharmaceutical Sciences and Faculty of Medicine, Vancouver, Canada (MS); School of Population Health, WHO Collaborating Centre for Climate Change and Health Impact Assessment, and enAble Institute, Faculty of Health Sciences, Curtin University, Perth, Australia (GP); Faculty of Medicine, Universitas Negeri Malang, Indonesia (GP); Saw Swee Hock School of Public Health, National University of Singapore, Singapore (WC)footnotetext: * Correspondence to Mohsen Sadatsafavi, Room 4110, Faculty of Pharmaceutical Sciences, 2405 Wesbrook Mall, Vancouver, British Columbia, V6T1Z3, Canada; email: mohsen.sadatsafavi@ubc.ca
 Background
 The underlying premise of reporting on the performance of biomarkers, tests, and prediction models (which we generally refer to as “markers”) is that performance metrics are transportable from one population to another.
-In this context, we define transportability as conditions under which certain indices of marker performance remain the same across populations.
-Examples of such indices include accuracy (sensitivity and specificity), predictive values, discrimination, and calibration.
 However, in practice, this premise more often than not fails to hold.
-Research has repeatedly shown the performance of markers can vary significantly across populations, especially when the prevalence of the outcome of interest differs1.
-For example, risk prediction models for cardiovascular disease have shown substantial degradation in performance when transported to a new population, leading to substantial risk of harm (defined as the net benefit of using the model being lower than not using it)2.
+Research has repeatedly shown the performance of markers can vary significantly across populations, especially when the prevalence of the outcome of interest differs4.
+For example, risk prediction models for cardiovascular disease have shown substantial degradation in performance when transported to a new population, leading to substantial risk of harm (defined as the net benefit of using the model being lower than not using it)5.
 The likelihood of harm was reduced when models were updated to account for the difference in prevalences between their source population and the target population.
-A similar phenomenon was observed in risk prediction models for acute exacerbations of obstructive lung disease3.
-In another example, the performance of machine learning models for detecting pneumonia on chest X-rays substantially declined in data from settings not used to train the model4.
-Often times, when transporting a marker for an outcome to a new population, some information about the characteristics of that population - colloquially referred to as case-mix- is available.
+In another example, the performance of machine learning models for detecting pneumonia on chest X-rays substantially declined in data from settings not used to train the model6.
+When transporting a marker for an outcome to a new population, some information about the characteristics of that population - generally referred to as case-mix- is often available.
 One of the most common information pieces is the outcome prevalence.
 For example, cancer registries often provide a good estimate of cancer risk in a population.
-This poses the challenge as to how such information can be used to revise our assessment of marker performance when transporting it to a new population.
-For risk predictions models that return a quantitative estimate of disease (or outcome) risk given patient characteristics, some form of model updating (or revision) is often needed to correct for the under-estimation or over-estimation of risks in the new population5.
-When the model is a logistic regression, the most basic form of such updating involves modifying the intercept of the model to account for difference in prevalence between the source and target populations1,6,7.
-The underlying justification for this approach is that differences in outcome prevalence are mostly due to factors that are distributed, and act, independently of predictors used in the model6.
-Because in a logistic regression model, changing the intercept is equivalent to applying an odds-ratio to predicted risks, this method can be generalized to applying a correcting odds-ratio to the outputs of any algorithm that predicts the risk of a binary outcome, including black-box (e.g., machine learning) models8.
+This poses the question as to how such information can be used to revise our assessment of marker performance.
+For risk predictions models that return a quantitative estimate of outcome risk given patient characteristics, some form of model revision is often needed to correct for the under-estimation or over-estimation of risks in the new population7.
+When the model is a logistic regression, the most basic form of such updating involves modifying the intercept of the model to account for difference in prevalence between the source and target populations1,8,9.
+Because in a logistic regression model, changing the intercept is equivalent to applying an odds-ratio to predicted risks, this method can be generalized to applying a correcting odds-ratio to the outputs of any risk prediction algorithm, including black-box (e.g., machine learning) models10.
 The transportability of predictive information across populations is an active area of research in predictive analytics and machine learning.
-A recent scoping review categorized methods aimed at developing transportable marker, or making an existing marker transportable to a new setting, based on whether they require access to data from the target population, and whether they are purely data-driven or require contextual knowledge about associations9.
-One common underlying framework in knowledge-driven approaches is causal graphs, particularly the pioneering work on selection diagrams by Pearl et al.10.
-Generally, these methods aim to identify and remove predictors whose association with the outcome varies across populations.
-More recent developments include the creation of graph surgery estimators that break the link between selection into the population and causal associations between the outcome and predictors whose association with the outcome is not consistent11,12.
-This approach has recently been used to study how common metrics of model performance for prognostic and diagnostic markers change across populations with difference case-mix13.
+A recent scoping review categorized methods aimed at developing transportable marker, or making an existing marker transportable to a new setting, based on whether they require access to data from the target population, and whether they are purely data-driven or require contextual knowledge about associations11.
+One common underlying framework in knowledge-driven approaches is causal graphs, particularly the pioneering work on selection diagrams by Pearl et al.12.
+Generally, these methods aim to identify and remove predictors whose association with the outcome varies across populations13,14.
+Causal diagrams have recently been used to study how common metrics of model performance for prognostic and diagnostic markers change across populations with difference case-mix15.
 Causal graphs are not the only model for causation.
-Another is the sufficient component causes (SCC) model14,15.
+Another is the sufficient component causes (SCC) model16,17.
 SCC is a foundational model in that it establishes fully deterministic relationships between causes and effects (rather than a representation of statistical dependencies as in causal graphs).
-This framework has recently been applied to explore the biologic plausibility of different link functions for modeling binary outcomes, resulting in proposals for more transportable measures of association such as the ‘switch relative risk’16 and more biology-aligned statistical models such as regression by composition17.
-In this paper, we use a simple SCC model to study the most basic prediction setup: a binary factor that is used as a marker for the risk of a binary outcome.
+This framework has recently been applied to explore the biologic plausibility of different link functions for modeling binary outcomes, resulting in proposals for more transportable measures of association18 and more biology-aligned statistical models19.
+In this paper, we use a parsimonious SCC model to study the most basic prediction setup: a binary factor that is used as a marker for the risk of a binary outcome.
 As a reference, we formulate this setup for prognostic markers, where adjustment for outcome prevalence seems to be a topical issue.
-We later discuss how symmetries can be used to investigate the transportability of diagnostic markers but leave deeper investigations to future work.
 This setup is the used to study marker transportability across populations, particularly as it relates to the variability in outcome prevalence.
@@
 A parsimonious SCC framework for prognostic markers
-The SCC framework assumes the existence of sets of sufficient causes that bring about an event18.
-Within each set, the causes are non-redundant (all elements with the set are required for the event to happen), but sets can act independently of each other19.
+The SCC framework assumes the existence of sets of sufficient causes that bring about an event20.
+Within each set, the causes are non-redundant (all elements with the set are required for the event to happen), but sets can act independently of each other22.
 Consider a binary prognostic marker, such as the presence or absence of BRCA1 mutation, and a binary outcome, such as the occurrence of breast cancer.
-While BRCA1 mutation is associated with increased cancer risk20, it is not a definitive marker: neither a negative BRCA1 mutation eliminates the risk of breast cancer, nor does its presence guarantee that cancer will occur.
+While BRCA1 mutation is associated with increased cancer risk23, it is not a definitive marker: neither a negative BRCA1 mutation eliminates the risk of breast cancer, nor does its presence guarantee that cancer will occur.
 The fact that neither positive nor negative marker values are definitive indicates that there are at least two other mechanisms at play.
@@
 On the other hand, cancer can also occur via pathways independent of the BRCA1 mutation, explaining why some breast cancer cases are BRCA1-negative.
-We now formalize a minimal setup that explains such indeterminism of a marker value for an outcome.
-We are interested in the properties of a binary prognostic marker $T$ for a binary outcome $D$.
+We now formalize a minimal causal setup.
+We consider a binary prognostic marker $T$ for a binary outcome $D$.
 For this marker to be informative but not definitive, at least two latent variables (or switches) must be present that can cause false-negative and false-positive responses.
-We can model these switches as follows:
+We model these switches as follows:
   • The latent binary variable $U$ represents universally required causes - for example factors that cause pre-cancers (in-situ cancer) to progress to malignancy.
@@
 We represent the contingency table by the sequence TP,FP,FN,TN{TP,FP,FN,TN}, where $TP=P(TD=1),FP=P(T\dot{D}=1),FN=P(\dot{T}D=1),TN=P(\dot{T}\dot{D}=1)$ are, respectively, true positive, false positive, false negative, and true negative probabilities.
-Our interest is in the four fundamental performance characteristics of binary markers: positive predictive value (PPV=P(D=1|T=1)=TP/(TP+FP){PPV}=P(D=1|T=1)=TP/(TP+FP)), negative predictive value (NPV=P(T=1|D=0)=TN/(FN+TN){NPV}=P(T=1|D=0)=TN/(FN+TN)), sensitivity (SE=P(T=1|D=1)=TP/(TP+FN){SE}=P(T=1|D=1)=TP/(TP+FN)), and specificity (SP=P(T=0|D=0)=TN/(FP+TN){SP}=P(T=0|D=0)=TN/(FP+TN)), as a function of prevalence ($P(D=1)=TP+FN$).
-By convention PPV and NPV are referred to as predictive values and SE and SP as accuracy metrics.
+Our interest is in the transportability of four fundamental performance characteristics of binary markers: positive predictive value (PPV=P(D=1|T=1)=TP/(TP+FP){PPV}=P(D=1|T=1)=TP/(TP+FP)), negative predictive value (NPV=P(T=1|D=0)=TN/(FN+TN){NPV}=P(T=1|D=0)=TN/(FN+TN)), sensitivity (SE=P(T=1|D=1)=TP/(TP+FN){SE}=P(T=1|D=1)=TP/(TP+FN)), and specificity (SP=P(T=0|D=0)=TN/(FP+TN){SP}=P(T=0|D=0)=TN/(FP+TN)), as a function of prevalence ($P(D=1)=TP+FN$).
+By convention PPV and NPV are referred to as predictive values, and SE and SP as accuracy metrics.
 Before proceeding, we note that one can create an equally parsimonious SCC setup by swapping the logical AND and OR in the above equation, resulting in the setup $D=(T\land V)\lor U$.
-However, this is mathematically symmetrical to our reference setup.
-This becomes obvious by considering that the complementary marker (whose positive and negative results are swapped) is a marker for not experiencing the outcome under the reference setup: $\dot{D}=(\dot{T}\lor\dot{V})\land\dot{U}$.
+However, this is mathematically symmetrical to our reference setup, as the complementary marker (whose positive and negative results are swapped) in this setup is a marker for not experiencing the outcome: $\dot{D}=(\dot{T}\lor\dot{V})\land\dot{U}$.
 Because of this, any pattern we observe for PPV in the main formulation is also observed for NPV in this alternative formulation, albeit in the opposite direction, and vice versa.
-Similarly, any pattern we observe for SE in the reference setup is also observable, in the opposite direction, for SP in this alternative setup.
-Table 1 shows the minimum SCC setup, and the resulting 2-by-2 contingency tables of marker and outcome values ($P(T,D)$), from which performance metrics can be derived.
+A similar symmetry exists between SE and SP in these setups.
+Table 1 shows these minimum SCC setups and the resulting contingency tables ($P(T,D)$).
 Table 1: Minimal configurations for a prognostic for the reference (left) and its symmetrical (right) setup, and the resulting contingency tables.
@@
 Mapping between contingency table and cause probabilities
-If all we know about marker performance in a populations is its contingency table, the full joint distribution $P(T,U,V)$ is unidentifiable, given its specification requires seven degrees of freedom but the contingency table only provides three.
-However, any joint distribution can be replaced by independent marginal distributions for $P(T=1)$, $P(U=1)$, and $P(V=1)$ that would result in the same contingency table.
+If all we know about marker performance in a populations is its contingency table, the full joint distribution of these causes among individuals in that population is unidentifiable, given its specification requires seven degrees of freedom, but the contingency table only offers three.
+However, any joint distribution can be replaced by independent marginal distributions for $P(T)$, $P(U)$, and $P(V)$ that would result in the same contingency table.
+Therefore, for studying contingency tables (and resulting marker performance metrics) for a single population, it is sufficient to model the probability of these causes independently across individuals within that population.
+However, the prevalence of these causes across populations can be correlated, and it is this dependence structure that determines the transportability of performance metrics.
+Within a population, given specifying three independent Bernoulli distributions for these probabilities requires three degrees of freedom, this setup is identifiable and has a unique (1:1) mapping to a given contingency table.
 We use the shorthand notations $P_{T}=P(T=1),P_{U}=P(U=1)$, and $P_{V}=P(V=1)$.
-Given specifying three Bernoulli distributions for these probabilities requires three degrees of freedom, this setup is identifiable and has a unique (1:1) mapping to a given contingency table.
-This 1:1 mapping can thus be formulated as follows:
+The 1:1 mapping can be formulated as follows:
   • From {TP,FP,FN,TNTP,FP,FN,TN} to $P_{T},P_{U},P_{V}$}:
-    $P_{T}=TP+FP,\quad P_{U}=TP/(TP+FP),\quad P_{V}=FN/[(1-P_{T})P_{U}]$
+    $P_{T}=TP+FP,\quad P_{U}=TP/P_{T},\quad P_{V}=FN/[(1-P_{T})P_{U}]$
   • From $P_{T},P_{U},P_{V}$} to {TP,FP,FN,TNTP,FP,FN,TN}:
@@
 Two common methods of transportability
-In this section we review the two conventional approaches for transporting a marker to a new population with potentially different outcome prevalence: transportation via accuracy (SE & SP), and transportation via predictive values (PPV & NPV).
 Ultimately, using a marker for decision-making entails interpreting its predictive values - i.e., $P(D|T)$.
@@
 Transportation by predictive values refers to the where predictive values from the source population are considered to be the same as in the target population.
-Given that $P(T,D)=P(T)P(D|T)$, this transportability means our estimate of the contingency table in the target population is determined by $P(T)$ in that population, combined with predictive values from the source population.
+Given that $P(T,D)=P(T)P(D|T)$, this transportability means the contingency table in the target population is determined by $P(T)$ in that population, combined with predictive values from the source population.
 This approach is more common for prognostic markers, a typical example being risk scoring tools that directly return an estimate of $P(D=1|T)$ for someone with marker value $x$.
@@
 Thus, under this transportability scheme we need access to prevalence in the target population to derive predictive values.
-This approach is the modus operandi for binary diagnostic markers, where the Bayes’ rule is used to combine pre-test probability with marker accuracy estimates to derive the post-test probability of the outcome21,22.
+This approach is the modus operandi for binary diagnostic markers, where the Bayes’ rule is used to combine pre-test probability with marker accuracy estimates to derive the post-test probability of the outcome24,25.
 Without any information that would distinguish the individual under evaluation, the pre-test probability is taken to be disease prevalence in the target population, which is equal to this mode of transportation.
 Returning to our running example, imagine we are transporting the above-mentioned prognostic marker to a population where $P_{T}=$ 0.250, $P_{U}=$ 0.750, $P_{V}=$ 0.500.
-The true contingency table and marker performance metrics are provided in Table 2.
+Outcome prevalence in this population is 0.622 - 32.7% higher than in the source population.
 Transportation by predictive values preserves PPV and NPV in the target population, but will result in SE and SP of 0.5 and 0.833.
-Now, imagine we know the target population prevalence, which is 0.622 (32.7% higher than in the source population).
+Now, imagine we know the target prevalence.
 If we consider transportation by accuracy, we will arrive at the PPV and NPV values of, respectively, 0.750 and 0.625.
-Prevalence adjustment for risk equations
-A common way that prognostic markers are transported across populations is in the format of a risk equation for $P(D|T)$, either explicitly, as is the case in regression-based prediction models, or implicitly, as in black-box (e.g., machine learning) models.
+The contingency table and marker performance metrics for the true population and those implied by the above-mentioned transportability methods are provided in Table 2.
+Prevalence-adjustment for risk equations
+A common way that prognostic markers are transported is in the format of a risk equation for $P(D|T)$, either explicitly, as is the case in regression-based prediction models, or implicitly, as in black-box (e.g., machine learning) models.
 A familiar modeling framework for binary outcomes is the logistic regression.
@@
 The first term on the right-hand side is the logit of prevalence, which is not a function of marker value.
-These derivations indicate that the practice of prevalence adjustment by the odds-ratio of prevalence between the source and target populations is equivalent to the application of Bayes’ theorem (where outcome prevalence is taken as the prior probability).
-This approach changes the locus of transportability from predictive values to the LR: for prevalence adjustment to make the marker transportable, the LR must remain constant across all levels of the marker.
+These derivations indicate that the practice of prevalence-adjustment by the odds-ratio of prevalence between the source and target populations is equivalent to the application of Bayes’ theorem.
+This approach changes the locus of transportability from predictive values to the LR.
 For binary markers, the LR is defined at two values.
@@
 For the marker to be transportable, both these LRs need to remain constant across populations, which will be the case if and only if both SE and SP remain constant.
-This in turn implies that prevalence adjustment for binary markers is equal to transportation by accuracy; i.e., considering SE and SP to be transportable.
 Under what conditions are performance metrics transportable?
-The above reasoning shows why adjustment for prevalence is generally expected to improve transportability, as the conventional wisdom is that SE and SP, being defined within the diseased and non-diseased groups, are less dependent on prevalence23.
+The above reasoning shows why adjustment for prevalence is generally expected to improve transportability, as the conventional wisdom is that SE and SP, being defined within the diseased and non-diseased groups, are less dependent on prevalence26.
 While this might be intuitive as a general observation, under the SCC framework, none of these metrics are truly intrinsic.
 Rather, they emerge as properties of the distribution of causes, and their transportability depends on how this distribution varies across populations.
-One can indeed create population-generating mechanisms where a given subset of these metrics remains constant while others vary.
-However, the SCC framework helps us examine the plausibility of such mechanisms.
-Table 3 shows conditions under which each of the four metrics remains constant (and therefore transportable) for a prognostic marker.
-Table 3: Conditions for Transportability of Metrics of Marker Performance
+One can create population-generating mechanisms where a given subset of these metrics remains constant while others vary.
+However, our parsimonious framework helps us examine the plausibility of such mechanisms.
+Table 3 shows conditions for the distribution of causes across populations under which each of the four metrics remains constant (and therefore transportable) for a prognostic marker.
+Table 3: Conditions for Transportability of Metrics of Marker Performance (c indicates a constant value)
 Metric Transportability condition
-PPV{PPV} $P(U=1|T=1)=c$
-NPV{NPV} $P(UV=1|T=0)=c$
-SE{SE} ${P(T=1|U=1)}/{P(\dot{T}V=1|U=1)}=c$
-SP{SP} ${P(T\dot{U}=1)}/[{P(T=0)-P(\dot{T}UV=1)}]=c$
+PPV{PPV} $P_{U}=c$
+NPV{NPV} $P_{U}P_{V}=c$
+SE{SE} $P_{V}(1-P_{T})/P_{T}=c$
+SP{SP} $P_{T}(1-P_{U})/[(1-P_{T})(1-P_{U}P_{V})]=c$
 SE: sensitivity; SP: specificity; PPV: positive predictive value; NPV: negative predictive value
-  • As an example of derivations, consider PPV. Its transportability means $P(D=1|T=1)=c$ (a constant). But $P(D=1|T=1)=(P(TU\dot{V}=1)+P(TUV=1))/P(T=1)=P(TU=1)/P(T=1)=P(U=1|T=1)$.
-For the PPV to be transportable, the prevalence of universal causes among marker-positive individuals should remain constant.
-For NPV, transportability requires that among test-negative individuals, the proportion in whom both the universal and alternative causes are present should remain constant.
-These conditions are both satisfied if the prevalences of $U$ and $V$ are stable across populations.
+  • As an example of derivations, consider PPV.
+Its transportability means $P(D=1|T=1)=c$ (a constant).
+But $P(D=1|T=1)=(P(TU\dot{V}=1)+P(TUV=1))/P(T=1)=P(TU=1)/P(T=1)=P(U=1|T=1)=P(U=1)$ (the last equality is based on our assumption of independence of cause distributions within a population).
+For the PPV to be transportable, the prevalence of universal causes should remain constant across populations.
+For NPV, transportability requires that the proportion in whom both the universal and alternative causes are present should remain constant.
+These conditions are both satisfied if $P_{U}$ and $P_{V}$ are stable across populations.
 In this scenario, the entirety of variation in outcome prevalence is attributable to variation in $T$.
 In our breast cancer example, this would mean populations vary in breast cancer prevalence only because they differ in the BRCA1 mutation rate.
-For many outcomes, including breast cancer, this assumption is unrealistic, as many factors contribute to the variability in breast cancer risk24.
+For many outcomes, this assumption is unrealistic (e.g., many factors contribute to the variability in breast cancer risk27).
 In comparison, conditions for the transportation by accuracy are more complicated.
-It does not seem possible to specify a simple model on how $P(T,U,V)$ should vary such that both SE and SP remain stable.
-Such transportability method for a prognostic marker, as is implied in prevalence adjustment, amounts to a very specific set of assumptions on how causes vary across populations.
+For SE, it requires the ratio of the probability of one cause ($V$) over the odds of another ($T$) to remain constant.
+For SP, it does not seem possible to specify a simple model on how causes should vary to ensure transportability.
+Overall, transportation by accuracy metrics for prognostic markers amounts to placing a very specific set of conditions on how causes vary across populations.
 How do performance metrics vary by prevalence under different population-generating mechanisms?
-In the SCC framework, the degree of transportability of a marker by a given performance metric depends on how the causes vary across populations.
+In our simple causal framework, the degree of transportability of a marker by a given performance metric depends on how the causes vary across populations.
 Despite the simple setup, complex patterns arise due to the non-linearities in the interplay among causes.
@@
 As explained above, under $T$-only variation, PPV and NPV remain transportable.
-This again highlights that transporting a marker by its predictive values (as in risk equations) means attributing change in prevalence entirely due to $T$, a very strong assumption in most real scenarios.
 Importantly, in all other conditions, PPV and NPV varied by prevalence.
@@
 Other forms of marker transportation under the SCC framework
-The SSC framework enables us to express our beliefs on how causal pathways vary across populations, resulting in algorithms that are transparent and explicit in their underlying assumptions.
+This framework enables us to express our beliefs on how causal pathways vary across populations, resulting in algorithms that are transparent and explicit in their underlying assumptions.
 Consider investigating a rare ancestral mutation as a prognostic marker for cancer that is known to have minimal variations across populations.
@@
 When transporting this marker, our focus would therefore be on scenarios where differences in $P(U,V)$ derive variations in cancer prevalence.
-The resulting transportation method will be different from transportation by predictive values or accuracy.
+The resulting transportation method will be different from transportation by predictive values or accuracy metrics.
 Indeed, both these transportation methods implicitly require that an increase in outcome prevalence be accompanied by change in marker positivity.
 Proportional-odds transportability method
+Methods based on causal graphs provide specific solutions for marker transportability if causal associations between specific factors that determine the marker and outcome values are known.
 But, what if the only information we have is differences in outcome prevalence across populations?
-We believe it is still worth visiting the plausibility of the assumptions under various transportability methods even under such a general case.
+Our parsimonious framework enables an overall assessment of the plausibility of the assumptions under various transportability methods even under such a general case.
 For transportability by predictive values, one can question as to why only one cause ($T$, the one we happen to be measuring) should be responsible for change in prevalence.
@@
 This “proportional odds” assumption is a new, distinct transportability approach that would generate different updated values of marker performance metrics compared with both conventional methods.
-Details of operationalizing this approach are provided in Appendix 2.
+Details of solving for this common odds-ratio are provided in the Appendix.
 In our numerical example (last row of Table 2), we need to apply an odds-ratio of 1.612 to the three causes to match the target population’s outcome prevalence.
-This odd-ratio can be derived via standard univariate root finding - given that there is a monotonical relationship between this odds-ratio and prevalence.
-It can also be expressed as the real root of a cubic equation for which standard solutions are available (see Appendix 2).
 The contingency table under this method of transportability will be {0.290, 0.060, 0.333, 0.318}.
@@
 The PPV and NPV are, respectively, 0.829 and 0.489.
-Formulating marker performance as a risk equation would thus result in $P(D|T)=0.171+0.317x$.
-Note that both the intercept and slope (coefficient for $T$) of the risk equation are updated, which is different from the risk equation under the assumption of transportability of predictive values (where both remain unchanged), or accuracy (where only the intercept is updated).
 Information loss under different transportability assumptions
@@
 We conducted brief simulation studies to explore such information loss under simple population-generating scenarios.
-The discrepancy between an assumed versus the true marker performance can be measured in different ways, each focusing on different consequences of such discrepancy (for example, through measures of discrimination, calibration, or prediction error).
-A more foundational approach is to measure information loss according to information theory.
-In this context, the relevant measure would be the Kullback-Leibler divergence ($D_{KL}$)25.
-This information-theoretic measure quantifies the discrepancy between a true distribution and a candidate distribution in terms of ‘excess surprise’, which can be interpreted as the additional number of bits required to encode information from the true distribution using the candidate distribution, rather than using the true distribution itself26.
+The discrepancy between an assumed versus the true marker performance can be measured in different ways, each focusing on certain aspects of performance (e.g., discrimination, calibration, or prediction error).
+A more foundational approach is to measure information loss by the Kullback-Leibler divergence ($D_{KL}$), an information-theoretic measure of the discrepancy between a true distribution and a candidate distribution28.
+$D_{KL}$ quantifies the additional number of bits required to encode information from the true distribution using the candidate distribution, rather than using the true distribution itself29.
 In our case, these distributions are the true contingency table ($P(T,D)$) in the target population versus the one implied by a given transportability algorithm.
 We modeled population-generating scenarios similar to those presented in Figure 1.
-Because in many realistic settings all three causes are likely to vary, we explored this setup with more depth.
-We modeled scenarios where changes in the three causes had different strengths of correlation with each other (ranging from changes being uncorrelated to fully correlated).
-We also modeled a ‘maximum entropy’ scenario where the three causes have independent uniform distributions; albeit unrealistic, this scenario represents maximum theoretical randomness across populations (therefore, marker performance is entirely untransportable)27.
-For each scenario, we simulated random pairs of source and target populations according to a given population-generating mechanism.
-When a cause was made variable, it was assumed that the logit-transformed probability of that cause had a standard Normal distribution (e.g., $\mbox{logit}(P_{T})\sim\mbox{Normal}(0,1)$ when where $P_{T}$ was variable across populations).
-For each pair, we applied the following three transportation strategies: by accuracy (SE and SP), by predictive values (PPV and NPV), and via the proportional-odds assumption explained previously.
-Divergence metrics were averaged across 10,000 simulations for each scenario.
+Because in many realistic settings all three causes are likely to vary, we explored this setup with more depth, covering scenarios were causes had positive correlation of varying strengths.
+We also modeled a ‘maximum entropy’ scenario where the three causes have independent uniform distributions; albeit unrealistic, this scenario represents maximum theoretical randomness, and thus places an upper limit on information loss30.
+For each scenario, we simulated 10,000 random pairs of source and target populations.
+For each pair, we applied the following transportation strategies: by accuracy, by predictive values, and via the proportional-odds assumption explained previously.
 Results are provided in Figure 2.
-This confirms that if the outcome prevalence variation is entirely due to $T$, PPV and NPV are fully transportable ($D_{KL}=0$).
+Overall, they show that no transportability method is universally better than others.
+They confirm that if the outcome prevalence variation is entirely due to $T$, PPV and NPV are fully transportable ($D_{KL}=0$).
 However, under other scenarios, transportation by predictive values resulted in information loss, sometimes substantially.
-Transportation by accuracy reduced information loss in some scenarios, but increased in others.
-This indicates that prevalence adjustment is not guaranteed to improve the transportability of a prognostic marker.
-In none of the scenarios was the loss zero, indicating that none of the modeled scenarios were compatible with stable SE and SP - again, due to the complex requirements for the stability of these metrics.
-The proportional-odds assumption also reduced information loss in some scenarios, but increased it in others.
-For the scenarios where the three causes changed together, transportation by accuracy generally reduced information loss compared with transportation by predictive values.
-In such scenarios, transportation via the proportional-odds assumption performed better than both conventional transportability methods.
+For transportability by accuracy, in none of the scenarios was the loss zero, indicating that none of the modeled scenarios were compatible with stable SE and SP - again, due to the complex requirements for the stability of these metrics.
+Notably, for the scenarios where the three causes changed together, the new, proportional-odds assumption performed better than both conventional transportability methods.
 Figure 2: Information loss associated with different methods for transporting a marker under various population-generating scenarios
@@
 Discussion
-We constructed a simple causal model under the SCC framework to study the transportability of prognostic markers that are partially informative about an outcome of interest but also generate false-positive and false-negative results.
-Several observations from this theoretical exploration deserve being highlighted.
-First, the common practice of transporting the marker by predictive values is questionable, as it relies on the strict assumption that the prevalences of causes other than those represented by the marker remain stable across populations.
-Further, we showed that the conventional prevalence adjustment method is the exact implementation of Bayes’ rule, thus changing the locus of transportability from predictive values to accuracy metrics (sensitivity and specificity).
+We constructed a parsimonious causal framework to study the transportability of prognostic markers.
+Several observations from this theoretical exploration deserve reiterating.
+First, the common practice of transporting prognostic markers by predictive values relies on the strong assumption that the prevalences of universal and alternative causes remain stable across populations.
+We showed that the conventional prevalence-adjustment method is the exact implementation of Bayes’ rule, thus changing the locus of transportability to accuracy (sensitivity and specificity).
 Nevertheless, no easily explainable or biologically plausible mechanism is likely to generate fully transportable accuracy metrics.
-In our explored scenarios regarding the distribution of causes across populations, transporting a prognostic marker by accuracy was not universally better than transporting by predictive values.
-However, in scenarios where causes were positively correlated, which is likely to be common, transporting by accuracy did reduce information loss compared with transportability by predictive values.
-However, in the same scenarios, the SSC-based ‘proportional odds’ method of transportation performed better.
-Our simulations were related to a selected scenarios and particular ways the causes vary across populations.
-Still, they should be sufficient to question some common assumptions and practices.
-These include the practice of advertising prognostic risk questions without emphasizing their dependence on prevalence, or considering sensitivity and specificity as intrinsic properties of markers.
+In our explored population-generating scenarios, transporting a prognostic marker by accuracy was not universally better than transporting by predictive values.
+However, in scenarios where causes were positively correlated, which is likely to be common, transporting by accuracy reduced information loss compared with transportability by predictive values.
+However, in the same scenarios, the new ‘proportional odds’ method of transportation performed better.
+Our numerical results were related to a selected ways the causes vary across populations.
+Still, they should be sufficient to question some common assumptions and practices, including the practice of advertising prognostic risk questions without emphasizing their dependence on case-mix and outcome prevalence, and considering sensitivity and specificity as intrinsic properties of markers.
 While these insights are from studying binary markers, the core findings can be extended to continuous markers and multi-variable risk equations.
 Transforming such markers ‘as is’ is equal to assuming that stratum-specific predictive values ($P(D|T)$) remain constant.
-This is equal to attributing variations in prevalence entirely to variations in $P_{T}$ (aka case-mix).
-Consider a multi-variable risk score such as the QRISK for cardiovascular diseases28.
-Even though this model includes up to 22 predictors, is it plausible that other causal pathways to cardiovascular diseases that might vary across populations, including environmental exposures, subtle lifestyle choices, access to and quality of preventive care, and yet-to-be-discovered genetic risk factors, are the same across populations?
-This assumption is needed for claiming that QRISK predictions, developed using primary care UK data, are transportable to other settings.
-On the other hand, conventional prevalence adjustment would indicate that stratum-specific likelihood ratio, i.e., $P(x|D=1)/P(x|D=0)$ would be transportable.
-This assumption will hold if the distribution of predicted risks among those who will experience the outcome and those who will not remains the same across diverse populations.
+This is equal to attributing variations in prevalence entirely to variations in $P_{T}$.
+Consider a multi-variable risk score such as the QRISK3 for cardiovascular diseases31.
+Even though this model includes up to 22 predictors, is it plausible that other causes of cardiovascular diseases, including environmental exposures, lifestyle choices, access to and quality of preventive care, and genetic risk factors, are the same across populations?
+This assumption is needed for claiming that QRISK3 predictions, developed using primary care UK data, are transportable to other settings.
+On the other hand, conventional prevalence-adjustment would indicate that stratum-specific likelihood ratios would be transportable (i.e., $P(T=x|D=1)/P(T=x|D=0)=c$ for all $x$).
+This assumption will hold if the distributions of predicted risks within outcome status strata remain the same.
 This condition places very specific constraints on how the causes of diseases should vary across populations.
 We focused on prognostic markers given the recent debates on transportability of prognostic information and the merit of prevalence-adjustment to improve transportability.
-The application of SCC framework for diagnostic markers is also important and deserves its own airing.
+The application of this framework for diagnostic markers is also important and deserves its own airing.
 For such markers, the path of causality if from the disease to the test.
 As such, an equivalent setup to our reference setup for a diagnostic marker would be $T=(D\lor V)\land U$.
-The conventional method of transporting diagnostic markers is by accuracy (e.g., the Bayes’ rule, which combines accuracy metrics from the source population with disease prevalence from the new population).
-From this setup, it is immediately obvious that transportation by accuracy requires that universal and alternative causes $U$ and $V$ remain stable across populations.
-This assumption is similar to the assumption of transportability by predictive values for prognostic markers.
-This strong requirement can explain why sensitivity and specificity tend to vary across populations29–31.
+From this setup, it is immediately obvious that transportation by accuracy (the default approach for diagnostic markers) requires that universal and alternative causes ($U$ and $V$) remain stable across populations.
+This strong requirement can explain why sensitivity and specificity tend to vary across populations2–4.
 How the SCC framework can inform transportability of diagnostic markers needs to be pursued in future work.
 How can these findings inform practice?
-If reliable, context-specific information is available on the distribution of specific predictors and their relationships with the outcome and with each other, methods based on causal graphs can be used to build transportable models or to design tailored transportation strategies11.
+we question the contemporary practice of advertising risk prediction models for outcome risk as transportable.
+If reliable, context-specific information is available on the distribution of predictors and their relationships with the outcome and with each other, methods based on causal graphs can be used to build transportable models or to design tailored transportation strategies13.
 In contrast, when only general information, such as outcome prevalence or test positivity rates, are available, the SCC framework can be used to formulate transportability algorithms that utilize such broad information.
-We question the contemporary practice of advertising risk prediction models for outcome risk as transportable.
-We believe risk equations should explicitly mention the outcome risk in their source population, and adjustment for prevalence should be advocated as a necessary step before using the equation in a population with different outcome prevalence.
-Yet, while conventional prevalence adjustment may serve as a reasonable choice, it cannot be assumed to consistently improve performance.
-Instead of fixed transportability rules, the SCC framework offers a foundation for updating marker performance in ways that are more grounded in mechanistic interactions among potential causes.
-This results in new transportability algorithms based on assumptions on how underlying causes vary, rather than keeping performance metrics, which are emergent properties of such causes, constant.
-An example is the proportional-odds prevalence adjustment method discussed above.
-It is motivated by the notion that oftentimes differences in outcome prevalence are multi-factorial.
-An agnostic stance that treats all causes as equally responsible is more defendable than blaming only one (as in transportation by predictive values) or assuming they move in unexplainable ways just to keep sensitivity and specificity constant (as in transportability by accuracy).
-Interestingly, in our numerical experiments, whenever conventional prevalence-adjustment performed better than transportability by predictive values, the SCC-based proportional-odds method performed better than both, indicating it might be a better choice.
-Of course, variations in this method are also possible.
-Imagine, for example, that in addition to knowing target prevalence, we have knowledge of marker positivity rate ($P_{T}$) in the new population (e.g., from an independent cross-sectional study).
-One can then directly adjust $P_{T}$ to match this value, and then apply a common odds-ratio to other causes to match target prevalence.
+Instead of fixed transportability rules, this framework offers a foundation for algorithms that reflect our assumptions on how underlying causes vary, rather than keeping performance metrics, which are emergent properties of such causes, constant.
 We conclude this paper by providing a few areas for further inquiry.
@@
 The issue of uncertain prevalence estimate is also applicable to conventional method.
-Empirical studies on conventional prevalence adjustment also estimated the prevalence from the sample that was subsequently used to assess marker performance - effectively taking its value as known2,3.
+Empirical studies on conventional prevalence adjustment also estimated the prevalence from the sample that was subsequently used to assess marker performance - effectively taking its value as known5,32.
 Further, our explorations were for when performance metrics are derived from a single source population.
@@
 This results in SCC-based model specifications for meta-analysis of marker performance studies, where the estimands are the parameters that govern the joint distribution of $T$, $U$, and $V$.
-This will provide an alternative to modeling the joint distribution of prevalence, sensitivity, and specificity (or prevalence and predictive values)32,33.
+This will provide an alternative to modeling the joint distribution of prevalence, sensitivity, and specificity (or prevalence and predictive values)33,34.
 The more explanatory nature of SCC-based specification might provide better fit to the data.
@@
 Our goal is to construct a predicted contingency table for the target population, defined by $TP^{*},FP^{*},FN^{*},TN^{*}$}.
-  1. 1.
-    Map from {TP,FP,FN,TNTP,FP,FN,TN} to $P_{T},P_{U},P_{V}$} (see text)
-  2. 2.
+Step 1. Map from {TP,FP,FN,TNTP,FP,FN,TN} to $P_{T},P_{U},P_{V}$} (see text)
+Step 2.
 Find the odds-ratio $x$ such that when applied to $P_{T}$, $P_{U}$, $P_{V}$, results in updated probabilities $P_{T}^{*}$, $P_{U}^{*}$, $P_{V}^{*}$ that correspond to the desired prevalence.
 Note that prevalence is equal to $P_{U}^{*}(P_{T}^{*}+P_{V}^{*}-P_{T}^{*}P_{V}^{*})$.
-Of note: $x$ can be found as the real root of the cubic equation $ax^{3}+bx^{2}+cx+d=0$ with
+Given that applying an odds-ratio $x$ to probability $p$ can be written as as the function $\frac{px}{1-p+px}$, we have
+$P_{T}^{*}(x)=\frac{P_{T}^{*}x}{1-P_{T}^{*}+P_{T}^{*}x}$ (similar for $P_{U}$ and $P_{V}$).
+This, we should solve for $x$ in $f(x)=\pi^{*}$ where
+$f(x)=P_{U}^{*}(x)[P_{T}^{*}(x)+P_{V}^{*}(x)-P_{T}^{*}(x)P_{V}^{*}(x)].$
+We consider non-degenerate scenarios where $0<P_{T}<1$, $0<P_{U}<1$, and $0<P_{V}<1$.
+Given that $f(x)$ is continuous, $f(0)=0$, and $\lim_{x\to\infty}f(x)=1$, there is always at least one real solution for $x$.
+Further, given that
+  • $\frac{dP_{T}^{*}(x)}{dx}>0$, $\frac{dP_{U}^{*}(x)}{dx}>0$, and $\frac{dP_{V}^{*}(x)}{dx}>0$, and
+  • $\frac{\partial f(x)}{\partial P_{T}^{*}(x)}>0$, $\frac{\partial f(x)}{\partial P_{U}^{*}(x)}>0$, $\frac{\partial f(x)}{\partial P_{V}^{*}(x)}>0$,
+$\frac{df(x)}{dx}>0$ for all $x$. As such, this solution is unique. This solution can be found using univariate root finding methods (e.g., uniroot() in R).
+Of note, re-arranging the terms reveals that $x$ can also be expressed as the real root of the cubic equation $ax^{3}+bx^{2}+cx+d=0$ with
 $a$ $=(1-\pi^{*})P_{T}P_{U}P_{V},$
@@
 $d$ $=-\pi^{*}(1-P_{T})(1-P_{U})(1-P_{V});$
-which can be solved using standard methods.
+which can be solved using standard methods (e.g., polyroot() in R).
   1. 3.